### v1

# DiMoDif: Discourse Modality-information Differentiation for Audio-visual Deepfake Detection and Localization

Abstract
Deepfake technology has rapidly advanced, posing significant threats to information integrity and societal trust. While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts but still altering the meaning, presents a more challenging detection scenario. We present a novel audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information. Our framework leverages features from deep networks that specialize in video and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery. To this end, DiMoDif employs a Transformer encoder-based architecture with a feature pyramid scheme and local attention, and optimizes the detection model through a composite loss function accounting for frame-level detections and fake intervals localization. DiMoDif outperforms the state-of-the-art on the Temporal Forgery Localization task by +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF. On the Deepfake Detection task, it outperforms the state-of-the-art by +30.5% AUC on AV-Deepfake1M, +2.8% AUC on FakeAVCeleb, and performs on-par on LAV-DF. Code available at https://github.com/mever-team/dimodif.
1 Introduction
Audio-visual deepfakes are AI-generated content involving manipulations to one or both modalities, sometimes at small time intervals, with the intention to deceive [45]. Deepfake content can widely spread online, and mislead viewers, contributing to the growing issue of information disorder. A common feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect manually. This necessitates the development of robust AI-based deepfake detection methods.
Deepfake detection is a growing area of research [54, 51, 46]. Existing approaches primarily focus on pixel-level analysis, which examines inconsistencies in gradient variations [65], color variations [30], or artifacts introduced during the generation process [11]. Also, feature-based methods analyze facial landmarks [72], temporal coherence [27], and other visual cues to identify inconsistencies in deepfake content [40]. Furthermore, there has been growing interest in detecting and localizing deepfakes in audio-visual content, with both tasks being particularly challenging due to the complex interplay between the two modalities. Key approaches include audio-visual synchronization that analyzes the consistency between visual and auditory cues namely face movements and speech [26, 9, 14], feature reconstruction learning with cross-reconstruction similarity estimation to detect anomalies prevalent in fake videos [76], and self-supervised approaches relying on the synchronization patterns learned from real videos only [28, 18].
Humans integrate multimodal sensory information, such as visual and auditory input, to extract meaningful features and perform a wide range of recognition tasks. Cognitive neuroscience research has extensively documented the interplay between video and audio [63, 64], often resulting in the alteration of perceptual content [70, 13] or even the induction of perceptual illusions [52, 60]. In the context of speech perception, the brain actively constructs content representations by combining visual, auditory, and contextual cues to predict ongoing and future utterances [12]. When visual information conflicts with auditory input, as in the McGurk effect [49] or in the case of deepfakes, increased prediction errors are observed, accompanied by significant changes in brain activity, which often manifest as higher-frequency neural oscillations and localized activation patterns, indicative of heightened cognitive effort [4].
(a) Example of speech recognition differences between lip-reading and speech-to-text models.
(b) Aggr. FakeAVCeleb [34].
(c) Aggr. LAV-DF [10].
(d) Aggr. AV-Deepfake1M [8].
Figure 1: The process of identifying audio-visual inconsistency for deepfake detection. In (a), a video’s visual and audio components are separately processed by lip-reading and speech-to-text models, then a difference score is calculated. In (b,c,d) the difference score distributions are illustrated in the form of violin plots for three different datasets.
We illustrate how state-of-the-art AI analysis tools could emulate this process in Figure 1. In Figure 1(a), an input video is first decomposed into its constituent visual and audio streams. These are then fed into state-of-the-art models for lip-reading and speech-recognition [44], and the resulting textual outputs are compared [^1] Python’s $\mathtt{difflib}$ is used (cf. https://docs.python.org/3/library/difflib.html).. Their difference is computed using the normalized Levenshtein distance, i.e. the ratio between the edit distance $\mathtt{delta}$ of the two sentences and the sum of their lengths, which results in scores between 0% (identical) and 100% (completely different). Figures 1(b), 1(c) and 1(d) illustrate aggregate results through violin plots computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. On average, lower difference is observed in real videos compared to fake ones across all datasets, indicating the suitability of this score as a measure of incongruity between audio and video. Even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which render naive thresholding on this score unsuitable for robust detection. This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, while samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
Motivated by the above, we propose Discourse Modality-information Differentiation (DiMoDif), a deep learning-based deepfake detection architecture that exploits the inter-modality differences in machine perception of speech. DiMoDif first decomposes the input video into its visual and audio streams, and extracts the respective video and audio representations based on state-of-the-art lip reading and speech-to-text pre-trained models [44]. At its core, DiMoDif employs a Transformer encoder architecture [66] with a feature pyramid scheme and local attention configuration, to spot frame-level cross-modal incongruities between the two modality-specific representations, and in that way detect and localize the deepfake forgeries. Note that low difference scores frequently occur in partly manipulated fake videos (e.g. Figure 1(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level. In addition, high difference scores occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf. Section 5.2). The architecture is optimized using a composite loss function that accounts for frame-specific detections, overlaps between predicted and ground truth fake intervals, and divergence in corresponding boundaries. We evaluate DiMoDif on Deepfake detection (DFD) and Temporal Forgery Localization (TFL) tasks using three audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8]. DiMoDif outperforms the state-of-the-art on FakeAVCeleb by +2.8% AUC, exhibits a significant performance increase of +30.5% AUC and +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF dataset. We also provide an analysis of DiMoDif’s generalization abilities, and an extensive ablation analysis to ground our methodological choices.
2 Related Work
2.1 Deepfake Detection Datasets
During the past years, several datasets have been proposed for video deepfake detection. Most of them focus solely on visual manipulations, thus including only the visual modality in their samples, e.g., DF-TIMIT [35], FaceForensics++ [58], DeeperForensics [32], Celeb-DF [41], WildDeepfake [81], and DF-Platter [48]. Our focus is on the detection of audio-visual deepfakes, which require content, manipulations and annotations in both modalities. Table 1 presents deepfake detection datasets that provide both visual and audio streams. DFDC [20] is a seminal large-scale benchmark in the field that contains approximately 128K samples. Although it provides audio-visual content and contains manipulations on both modalities, (i) it does not provide separate manipulation annotations per modality, (ii) fake audio parts appear in only 4% of the total number of videos, while fake visual parts in 84% [^2] Based on a later estimation: https://www.kaggle.com/datasets/basharallabadi/dfdc-video-audio-labels., and (iii) it contains videos where the depicted individual does not talk or the speaking person is not depicted and the depicted one does not talk. These features render DFDC inappropriate for audio-visual deepfake detection. KoDF [36] is another large-scale benchmark containing 237K videos of Korean subjects. Although it provides both the visual and audio components of its samples, it contains only visual manipulations, which makes it unsuitable for our problem. FakeAVCeleb [34], contains 21K manipulated and 0.5K real videos with manipulations on both modalities and corresponding annotations. Due to its size it is commonly used as an evaluation benchmark; here, we consider it for training and evaluation purposes with real video additions from the VoxCeleb2 dataset [15]. LAV-DF [10] is a recently introduced dataset that contains 136K real and fake videos with manipulations and targets on both modalities for two tasks, Deepfake Detection (DFD) and Temporal Forgery Localization (TFL). Finally, AV-Deepfake1M [8] is the largest audio-visual deepfake detection benchmark to date containing over 1.1M real and fake videos with manipulations and targets on both modalities, for both the DFD and TFL tasks. In our experimental analysis, we consider FakeAVCeleb, LAV-DF, and AV-Deepfake1M datasets, which meet all requirements of the audio-visual deepfake detection task. We also consider DFDC and KoDF for cross-dataset generalization evaluation experiments.
content manip. target
dataset size A V A V A V joint task
DFDC [20] 128,154 DFD
KoDF [36] 237,942 DFD
FakeAVCeleb [34] 21,544 DFD
LAV-DF [10] 136,304 DFD,TFL
AV-Deepfake1M [8] 1,146,760 DFD,TFL
Table 1: Deepfake detection datasets providing both video and audio. Size determines the number of samples, content specifies the provided modalities, manipulation specifies which modality has been manipulated, target specifies for which modality the dataset provides ground truth, while task is reported in the last column.
2.2 Deepfake Detection Approaches
Many deepfake detection approaches consider only video-level manipulations and processing [62, 58, 78, 38, 1, 17, 55]. Although these are only partly related to DiMoDif  we compare our method with them following previous practice. Recent works in deepfake detection have focused on leveraging both visual and audio information through learning mechanisms based on multimodal fusion, contrastive learning, and self-supervised-learning, aiming to uncover cross-modal inconsistencies in deepfake videos. Specifically, [80] demonstrated the effectiveness of an attention-based detection framework that identifies the violation of cross-modal synchronization patterns in deepfake videos, while [71] proposed a framework exploiting audio-visual inconsistency by encoding spatio-temporal information, fusing multimodal features, and then classifying based on inter- and intra-modal disharmony levels. Similarly, [56, 31, 82] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms. UMMAFormer [76] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives. The contrastive learning approach is also adopted by other works for comparing inter-modal representations within a video [14] or between different videos [77]. Other papers focus on feature similarity to identify audio-visual inconsistencies [26], and [59] compares features of the video’s lip sequence with a lip sequence generated from the audio input. Self-supervised approaches that leverage intrinsic synchronization of real videos have also been utilized [75, 25, 28, 22, 18, 53, 73, 21] to tackle the generalization challenge faced by methods trained on specific deepfake generation models. A similar motivation with ours is presented in [7, 39], which rely on a simple threshold on the divergence between visual and audio speech recognition model outputs. However, as shown in Figure 1 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M. In contrast, our approach extracts lip-reading and speech-to-text content representations that are then processed by a Transformer-based fusion mechanism that is capable of identifying discriminative speech features on top of speech divergence ones. Pre-trained lip-reading models have also been utilized as feature extractors in [29], but for visual-only deepfake detection. In addition, [57] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model. Also, specialized representation learning has been explored by [47] that considers modality emotion consistency and [69] that considers articulatory and lip movement consistency. Finally, several works [42, 50, 61, 74, 5] of the Temporal Action Localization (TAL) domain are considered as relevant by the deepfake detection literature.
3 Methodology
3.1 Problem Formulation
Consider a dataset $\mathfrak{D}=\{\mathfrak{v}_{i},y_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ where $\mathfrak{v}_{i}=\{v,a\}$ denotes a video, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^3] Utterances mainly comprise monophonic signals, besides conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples. $y_{i}\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}_{i}=\{y^{v}_{i,1},\dots,y^{v}_{i,f},y^{a}_{i,1},\dots,y^{a}_{i,f}\}$, with $y^{m}_{i,j}=(d^{m,s}_{i,j},d^{m,e}_{i,j},a^{m}_{i,j})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{i,j}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $j$, along with the corresponding distances $d^{m,s}_{i,j},d^{m,e}_{i,j}\in\mathbb{R}$ between the frame $j$ and the onset $s^{m}_{i,j}$ and offset $e^{m}_{i,j}$ of the fake interval, if the frame $j$ is fake. We train deep networks $\mathcal{F}$, the architecture of which is analysed in Section 3.2, on the DFD task $\hat{y}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ and the TFL task $\hat{\mathbf{y}}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ separately, while the predictions $\hat{\mathbf{y}}_{i}$ are decoded by:
$\hat{a}^{m}_{i,j}>0.5\text{, }\hat{s}^{m}_{i,j}=j-\hat{d}^{m,s}_{i,j}\text{, }\hat{e}^{m}_{i,j}=j-\hat{d}^{m,e}_{i,j}$ (1)
$\hat{\cdot}$ denotes prediction while its absence ground truth. Henceforth, we omit sample index $i$ to reduce notation complexity.
3.2 Architecture
Our architecture is illustrated in Figure 2. From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the lip-reading and speech-to-text models proposed in [44]. Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected. For that reason, the proposed architecture that identifies such patterns is named Discourse Modality-information Differentiation (DiMoDif).
The discourse features, $\mathbf{v}$ and $\mathbf{a}$, are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers. Then, we add trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ and employ a separation token $\mathbf{s}\in\mathbb{R}^{d}$ to encode modality information. An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{j}\in\mathbb{R}^{d}$, with $j=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [66]. $e$ depends on the task, being equal to 2 in DFD experiments and equal to 1 in TFL experiments where $\mathbf{c}$ is absent. Then, we extract $d$-dimensional feature pyramids corresponding to all $2\cdot f$ visual and audio tokens deriving from all $l$ layers of the Transformer encoder, denoted by $\mathbb{R}^{(2\cdot f)\times l\times d}\ni\mathbf{z}\subset\mathcal{T}(\mathbf{t})$. To do so, $\mathcal{T}$ processes $\mathbf{t}$ as:
$\tilde{\mathbf{z}}_{j}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{j-1}))$ (2)
$\mathbf{z}_{j}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{j}))+\tilde{\mathbf{z}}_{j}$ (3)
where $j=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [66] with $r$ number of heads, LN denotes Layer Normalization [37], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$. Then, $\mathbf{z}=\oplus\{\mathbf{z}_{j}\}_{j=1}^{l}$, omitting the tokens that correspond to $\mathbf{c}$ and $\mathbf{s}$. Given that short-term inter-modality inconsistency detection is key to our tasks, local attention is preferable in contrast to global; besides, it is computationally more efficient. Thus, we limit the Transformer’s attention to a local window $[j-q/2,j+q/2]$ of $q$ tokens around each token $j$.
Figure 2: The DiMoDif architecture.
Finally, in DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\mathbf{z}$, denoted by $\mathbf{z}^{\prime}\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
$\mathbf{z}^{\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime}\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
$\mathbf{z}^{\prime\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime\prime}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
$\hat{y}$ $=\mathbf{z}^{\prime\prime\prime}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
[…]
3.3 Objective Function
For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective. For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals. Detailed mathematical expressions explaining the computation of both loss functions can be found in the Appendix A.
4 Experimental Setup
4.1 Datasets
We consider FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8] datasets for training and evaluation. Appendix B provides details on class and split sizes. FakeAVCeleb contains 21K fake and 500 real samples, thus for balance we added 20K more from VoxCeleb2 dataset [15] including 10K of male and 10K of female speakers of several ethnicities talking in English [^4] Language identification conducted with https://github.com/speechbrain/speechbrain., reaching a total of 41,544 samples randomly split based on a 80-10-10 identity separation. LAV-DF has been released with standard splits of 78K training, 31K validation, and 26K test samples. AV-Deepfake1M has 746K training, 57K validation, and 343K test samples the metadata of which have not been released. We also evaluate DiMoDif on the test set of DFDC [20] and a randomly selected 6K-sized sample from KoDF [36] proportionally spanning all synthesis methods and real samples, to assess generalization.
4.2 Evaluation
For DFD, we use accuracy (ACC), Average Precision (AP), and area under ROC curve (AUC). For TFL, we use Average Precision at $p$ (AP@$p$), and Average Recall at $n$ (AR@$n$). We report metrics per dataset in accordance to previous works for comparability. On AV-Deepfake1M we evaluate through Codabench [^5] https://deepfakes1m.github.io/evaluation. Competing methods are audio-visual deepfake detectors, face manipulation detectors, and temporal action localization models, discussed in Section 2.
4.3 Implementation Details
We train DiMoDif for 100 epochs with early stopping patience 10 and checkpoint based on validation metrics, AUC for DFD and the sum of AP@$p$ and AR@$n$ for TFL. Batch size is set to 64, initial learning rate to 0.001 with ReduceLROnPlateau scheduler [^6] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html and Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$. Also, we set a maximum sequence length $f$ to 600 and zero-pad the smaller. We consider a grid for $\mathcal{T}$’s size, with $(d,r,u)\in\{(32,2,1),(64,2,1),(64,4,1),(128,4,2),(256,8,2)\}$ and number of layers $l\in\{1,3,5\}$. Grid search for AV-Deepfake1M is performed with 200K training samples to reduce resource requirements, but retrain the best configuration on the whole set. We spent $\sim$3900 GPU hours for training on NVIDIA GeForce RTX 3090 Ti GPUs.
5 Results
5.1 Temporal Forgery Localization
Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
MDS [14] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
AGT [50] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
BMN [42] $\mathcal{V}$ 24.0 7.6 0.1 53.3 41.2 31.6 26.9
BMN (I3D) [42] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
ActionFormer [74] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
UMMAFormer [76] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
Table 2: Temporal forgery localization results on LAV-DF [10]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Method Modality AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@50 AR@30 AR@20 AR@10 AR@5
Meso4 [1] $\mathcal{V}$ 09.86 06.05 02.22 00.59 38.92 38.91 38.81 36.47 26.91
MesoInception4 [1] $\mathcal{V}$ 08.50 05.16 01.89 00.50 39.27 39.22 39.00 35.78 24.59
EfficientViT [17] $\mathcal{V}$ 14.71 02.42 00.13 00.01 27.04 26.99 26.43 23.90 20.31
TriDet+VideoMAEv2 [61, 67] $\mathcal{V}$ 21.67 05.83 00.54 00.06 20.27 20.23 20.12 19.50 18.18
TriDet+InternVideo [61, 68] $\mathcal{V}$ 29.66 09.02 00.79 00.09 24.08 24.06 23.96 23.50 22.55
ActionFormer+VideoMAEv2 [74, 67] $\mathcal{V}$ 20.24 05.73 00.57 00.07 19.97 19.93 19.81 19.11 17.80
ActionFormer+InternVideo [74, 68] $\mathcal{V}$ 36.08 12.01 01.23 00.16 27.11 27.08 27.00 26.60 25.80
BA-TFD [10] $\mathcal{AV}$ 37.37 6.34 0.19 0.02 45.55 40.37 35.95 30.66 26.82
BA-TFD+ [9] $\mathcal{AV}$ 44.42 13.64 0.48 0.03 48.86 44.51 40.37 34.67 29.88
UMMAFormer [76] $\mathcal{AV}$ 51.64 28.07 7.65 1.58 44.07 43.93 43.45 42.09 40.27
DiMoDif (ours) $\mathcal{AV}$ 86.93 75.95 28.72 5.43 81.57 80.85 80.25 78.84 76.64
Table 3: Temporal forgery localization results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Tables 2 and 3 present the performance of our method in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8] datasets, respectively. On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50,20}, is second to best wrt AR@10, and third wrt AP@{0.5,0.75,0.95}. UMMAFormer [76], BA-TFD+ [9] and ActionFormer [74] exhibit similar performance, while the rest of the models have lower localization abilities. On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase. The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
Method Modality AUC
F3-Net [55] $\mathcal{V}$ 52.0
MDS [14] $\mathcal{AV}$ 82.8
EfficientViT [17] $\mathcal{V}$ 96.5
BA-TFD [10] $\mathcal{AV}$ 99.0
UMMAFormer [76] $\mathcal{AV}$ 99.8
DiMoDif (ours) $\mathcal{AV}$ 99.6
Table 4: Deepfake detection results on LAV-DF [10]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance.
Method Modality AUC ACC
Video-LLaMA (13B) E5 [75] $\mathcal{AV}$ 50.7 25.1
LipForensics [29] $\mathcal{V}$ 51.6 68.8
Face X-Ray [38] $\mathcal{V}$ 61.5 73.8
Meso4 [1] $\mathcal{V}$ 50.2 75.0
MesoInception4 [1] $\mathcal{V}$ 50.1 75.0
SBI [62] $\mathcal{V}$ 65.8 69.0
MDS [14] $\mathcal{AV}$ 56.6 59.4
DiMoDif (ours) $\mathcal{AV}$ 96.3 83.3*
Table 5: Deepfake detection results on AV-Deepfake1M [8]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. *Validation set accuracy is reported as Codabench only provides AUC score.
Category
Method Modality Pre-training RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
Unsupervised AVBYOL [25] $\mathcal{AV}$ LRW [16] 50.0 50.0 73.4 61.3 88.7 80.8 60.2 33.8 73.2 61.0 73.9 59.2
VQ-GAN [21] $\mathcal{V}$ LRS2 [2] - - 50.3 49.3 57.5 53.0 49.6 48.0 62.4 56.9 55.0 51.8
AVAD [22] $\mathcal{AV}$ LRS2 [2] 62.4 71.6 93.6 93.7 95.3 95.8 94.1 94.3 93.8 94.1 94.2 94.5
AVAD [22] $\mathcal{AV}$ LRS3 [3] 70.7 80.5 91.1 93.0 91.0 92.3 91.6 92.7 91.4 93.1 91.3 92.8
Supervised Xception [58] $\mathcal{V}$ ImageNet [19] - - 88.2 88.3 92.3 93.5 67.6 68.5 91.0 91.0 84.8 85.3
LipForensics [29] $\mathcal{V}$ LRW [16] - - 97.8 97.7 99.9 99.9 61.5 68.1 98.6 98.7 89.4 91.1
AD DFD [80] $\mathcal{AV}$ Kinetics [33] 74.9 73.3 97.0 97.4 99.6 99.7 58.4 55.4 100. 100. 88.8 88.1
FTCN [78] $\mathcal{V}$ 96.2 97.4 100. 100. 77.4 78.3 95.6 96.5 92.3 93.1
RealForensics [28] $\mathcal{V}$ LRW [16] - - 88.8 93.0 99.3 99.1 99.8 99.8 93.4 96.7 95.3 97.1
DiMoDif (ours) $\mathcal{AV}$ 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
Table 6: Deepfake detection results on FakeAVCeleb [34]. Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio. Bold indicates best and underline second to best performance. Supervised methods that use pre-training are fine-tuned on FakeAVCeleb, while unsupervised methods are not trained with labels and fake examples.
Test dataset
FakeAVCeleb [34] LAV-DF [10] AV-Deepfake1M [8] DFDC [20] KoDF [36]
ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC
Training dataset FakeAVCeleb [34] 98.1 99.9 99.9 39.3 59.5 61.9 25.3 55.1 55.1 51.4 62.7 60.8 65.5 77.4 32.7
LAV-DF [10] 43.6 70.7 71.3 95.6 99.6 99.6 32.1 69.7 66.2 51.8 52.9 54.5 43.8 88.8 66.5
AV-Deepfake1M [8] 56.3 82.0 85.4 59.2 83.2 81.5 87.7 97.3 97.3 51.3 50.5 49.7 77.0 77.7 33.4
Table 7: Generalization performance on Deepfake detection task.
AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@100 AR@50 AR@30 AR@20 AR@10 AR@5
Test dataset: LAV-DF
Training dataset LAV-DF [10] 95.5 87.9 54.6 20.6 94.2 93.7 93.2 92.7 91.4 89.2
AV-Deepfake1M [8] 59.8 26.2 2.3 0.2 75.3 72.3 69.7 67.3 62.5 56.2
Test dataset: AV-Deepfake1M
LAV-DF [10] 16.8 4.6 0.3 0.0 31.9 30.1 28.6 27.0 23.6 19.9
AV-Deepfake1M [8] 96.0 86.2 34.5 6.9 89.2 88.7 88.2 87.7 86.4 84.3
Table 8: Generalization performance on Temporal Forgery Localization task.
5.2 Deepfake Detection
Tables 5, 5 and 6, present comparative results in terms of video-level Deepfake Detection (DFD) on LAV-DF [10], AV-Deepfake1M [8], and FakeAVCeleb [34], respectively. On LAV-DF, DiMoDif performs on-par with the state-of-the-art method UMMAFormer [76], achieving 99.6 AUC, better than the remaining competitive methods. On AV-Deepfake1M dataset, DiMoDif outperforms all state-of-the-art competitive methods by a significant absolute performance increase of +30.5% in terms of AUC [^7] Accuracy (ACC) is also reported by other papers; however, Codabench does not provide it for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase.. On FakeAVCeleb, DiMoDif outperforms all competitive methods on average achieving +4.1% absolute AP increase, and +2.8% absolute AUC increase, as well as in most forgery types. Finally, challenging real samples with normalized Levenshtein distance over 0.8 (cf. Section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt a naive distance-thresholding approach.
5.3 Generalization
This paper is the first to conduct generalization experiments on the used datasets. Table 7 presents DiMoDif’s cross-dataset performance on DFD. Training on FakeAVCeleb [34] leads to poor performance on LAV-DF, AV-Deepfake1M, and DFDC but to moderate performance on KoDF with 77.4 AP. Training on LAV-DF [10] leads to moderate performance on FakeAVCeleb and AV-Deepfake1M, low performance on DFDC but good performance on KoDF with 88.8 AP. Training on AV-Deepfake1M [8] dataset leads to high scores on FakeAVCeleb and LAV-DF with 85.4 AUC and 81.5 AUC respectively, moderate performance on KoDF with 77.7 AP, but low performance on DFDC. In general, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.1. Our models focus on cross-modal information differences, while many DFDC videos do not contain talking or the talking person is different from the depicted one; thus, a visual-only face manipulation detector would likely lead to better results. Appendix C provides detailed generalization evaluation on FakeAVCeleb with scores per forgery type. Table 8 presents cross-dataset performance on TFL. Training on AV-Deepfake1M leads to acceptable performance on LAV-DF, in contrast to training on LAV-DF and evaluating on AV-Deepfake1M.
5.4 Ablation Analysis
(a) Window size $q$ (LAV-DF)
(b) Window size $q$ (AV-Deepfake1M)
(c) Feature pyramid use (LAV-DF)
(d) Feature pyramid use (AV-Deepfake1M)
(e) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
(f) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
Figure 3: Ablation and hyperparameter tuning analysis.
Figure 3 presents an ablation analysis that supports our methodological choices. For AV-Deepfake1M we report validation performance as conducting the ablations through Codabench would be impractical. These results pertain to TFL, while further ablations as well as the same analysis wrt DFD can be found in Appendix D. Figures 3(a) and 3(b) indicate that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$. Also, the feature pyramid scheme provides a small performance increase as indicated by Figures 3(c) and 3(d). The size of $\mathcal{T}$ plays crucial role in achieving maximum performance as shown in Figures 3(e) and 3(f), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$. Specifically, the larger the model the higher its performance.
6 Conclusions
In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech. It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities. It considers a video and audio speech recognition feature extraction stage, and a Transformer encoder-based learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries. An extensive evaluation study indicates the effectiveness of our framework reaching state-of-the-art performance and in several cases surpassing it. Our model is limited by the languages recognized by the used speech-recognition backbones, although it exhibits good generalization to Korean. Also, it assumes one continually visible speaking person, at least moderate video quality and face size to detect face landmarks, and requires both modalities for manipulation detection. Future work includes joint training on Deepfake Detection and Temporal Forgery Localization tasks to assess potential benefits and extension to multiple speakers.
7 Acknowledgments
This work has been partially funded by the Horizon Europe projects AI4Trust (GA No. 101070190) and vera.ai (GA No. 101070093).
Appendix A Objective function details
For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective:
$\mathfrak{L}_{DFD}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{j=1}^{l}y[m]\text{log}\hat{y}[j,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[j,m])$ (7)
where $j$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
For the TFL task, we consider a combination of three loss functions. For clarity we omit to also include the symbols for the averaging across $l$ layers, which however takes place as well. Specifically, we use:
  1. 1.
    The focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent among each video’s frames:
    $\mathfrak{L}_{focal}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}-\alpha_{t}^{m,j}(1-p_{t}^{m,j})^{\gamma}\text{log}(p_{t}^{m,j})$ (8)
    in which:
    $\alpha_{t}^{m,j}=\begin{cases}\alpha&\text{if }y^{m}_{j}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
    $p_{t}^{m,j}=\begin{cases}\hat{y}^{m}_{j}&\text{if }y^{m}_{j}=1\\ 1-\hat{y}^{m}_{j}&\text{otherwise}\end{cases}$ (10)
    while $\alpha$ and $\gamma$ are hyperparameters.
  2. 2.
    The DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals:
    $\mathfrak{L}_{DIoU}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{j},\mathbf{b}^{m}_{j})+\\ \frac{\rho^{2}(\hat{b}^{m}_{j},b^{m}_{j})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,j}$ (11)
    where $j$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{j}=[s^{m}_{j},e^{m}_{j}]$ denotes a time interval, $b^{m}_{j}=0.5\cdot(e^{m}_{j}+s^{m}_{j})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake.
  3. 3.
    The smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals:
    $\mathfrak{L}_{1}^{smooth}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{j}-\hat{s}^{m}_{j})\\ +h(e^{m}_{j}-\hat{e}^{m}_{j})\bigg{)}\cdot\mathds{1}_{m,j}$ (12)
    in which $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake, and $h$ is defined as:
    $h(x)=\begin{cases}0.5x^{2}&\text{if }\lvert x\rvert<1\\ \lvert x\rvert-0.5&\text{otherwise}\end{cases}$ (13)
We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\mathds{1}_{m,j}$:
$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{focal}+\mathfrak{L}_{DIoU}+\mathfrak{L}_{1}^{smooth}\big{)}/p$ (14)
We finally average both $\mathfrak{L}_{DFD}$ and $\mathfrak{L}_{TFL}$ across batch samples during training.
Appendix B Dataset class and split sizes
Table 9 provides details on dataset class and split sizes. Although, FakeAVCeleb paper [34] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples instead of the commonly reported 20,000 size for this dataset. Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”. Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
FakeAVCeleb LAV-DF AV-Deepfake1M
train val test total train val test total train val test total
FVFA 8,678 1,052 1,105 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
FVRA 7,760 975 974 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
RVFA 399 49 52 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
RVRA 17,766 1,535 1,199 20,500 21,254 8,271 6,906 36,431 186,666 14,235 - -
total 34,603 3,611 3,330 41,544 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
Table 9: Dataset class and split sizes. F: fake, R: real, V: video, A: audio.
Appendix C Generalization on FakeAVCeleb
Table 10 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M especially in some cases such as FVFA-WL and FVFA-FS reaching 95.6 AUC and 94.5 AUC, respectively. Training on LAV-DF also results in good generalization especially in FVFA-FS case with 90.0 AUC.
Category
Training dataset RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
FakeAVCeleb [34] 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
LAV-DF [10] 40.2 74.3 77.1 87.2 78.2 87.8 72.1 90.0 70.0 86.9 74.4 88.0
AV-Deepfake1M [8] 18.8 80.8 63.5 80.3 85.6 95.6 74.2 94.5 75.9 93.8 74.8 91.1
Table 10: Detailed generalization performance for each forgery type of the FakeAVCeleb dataset.
Appendix D Ablation Analysis (further experiments)
Figures 4(a) and 4(b) indicate little to no sensitivity to the focal loss’ hyperparameter $\alpha$. Figures 4(c) and 4(d) indicate that the use of a learning rate scheduler is beneficial however the type of scheduler is not significantly affecting the results.
(a) Focal loss hyperparmeter $\alpha$ (LAV-DF)
(b) Focal loss hyperparmeter $\alpha$ (AV-Deepfake1M)
(c) Learning rate scheduler (LAV-DF)
(d) Learning rate scheduler (AV-Deepfake1M)
Figure 4: Ablation analysis on Temporal Forgery Localization task.
Figure 5 indicates that on DFD window size $q=15$ is again optimal, but performance is not increased with feature pyramids nor with different learning rate schedules. Also, medium-sized models achieve maximum performance.
(a) Window size $q$ (FakeAVCeleb)
(b) Window size $q$ (LAV-DF)
(c) Window size $q$ (AV-Deepfake1M)
(d) Feature pyramid use (FakeAVCeleb)
(e) Feature pyramid use (LAV-DF)
(f) Feature pyramid use (AV-Deepfake1M)
(g) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (FakeAVCeleb)
(h) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
(i) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
(j) Learning rate scheduler (FakeAVCeleb)
(k) Learning rate scheduler (LAV-DF)
(l) Learning rate scheduler (AV-Deepfake1M)
Figure 5: Ablation and hyperparameter tuning analysis on Deepfake Detection task.

### v2

@@
 Abstract
-Deepfake technology has rapidly advanced, posing significant threats to information integrity and societal trust.
-While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts but still altering the meaning, presents a more challenging detection scenario.
-We present a novel audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information.
-Our framework leverages features from deep networks that specialize in video and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery.
-To this end, DiMoDif employs a Transformer encoder-based architecture with a feature pyramid scheme and local attention, and optimizes the detection model through a composite loss function accounting for frame-level detections and fake intervals localization.
-DiMoDif outperforms the state-of-the-art on the Temporal Forgery Localization task by +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF.
-On the Deepfake Detection task, it outperforms the state-of-the-art by +30.5% AUC on AV-Deepfake1M, +2.8% AUC on FakeAVCeleb, and performs on-par on LAV-DF.
+Deepfake technology has rapidly advanced and poses significant threats to information integrity and trust in online multimedia.
+While significant progress has been made in detecting deepfakes, the simultaneous manipulation of audio and visual modalities, sometimes at small parts or in subtle ways, presents highly challenging detection scenarios.
+To address these challenges, we present DiMoDif (fig.
+1), an audio-visual deepfake detection framework that leverages the inter-modality differences in machine perception of speech, based on the assumption that in real samples – in contrast to deepfakes – visual and audio signals coincide in terms of information.
+DiMoDif leverages features from deep networks that specialize in visual and audio speech recognition to spot frame-level cross-modal incongruities, and in that way to temporally localize the deepfake forgery.
+To this end, we devise a hierarchical cross-modal fusion network, integrating adaptive temporal alignment modules and a learned discrepancy mapping layer to explicitly model the subtle differences between visual and audio representations.
+Then, the detection model is optimized through a composite loss function accounting for frame-level detections and fake intervals localization.
+DiMoDif outperforms the state-of-the-art on the Deepfake Detection task by 30.5 AUC on the highly challenging AV-Deepfake1M, while it performs exceptionally on FakeAVCeleb and LAV-DF.
+On the Temporal Forgery Localization task, it outperforms the state-of-the-art by 47.88 AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF.
 Code available at https://github.com/mever-team/dimodif.
 1 Introduction
-Audio-visual deepfakes are AI-generated content involving manipulations to one or both modalities, sometimes at small time intervals, with the intention to deceive [45].
-Deepfake content can widely spread online, and mislead viewers, contributing to the growing issue of information disorder.
-A common feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect manually.
+Figure 1: Partial audio-visual manipulation leads to different visual and audio speech predictions. DiMoDif detects and localizes the fake part based on feature space incongruity.
+Audio-visual deepfakes are AI-generated multimedia involving manipulations to one or both modalities, sometimes in small parts, with the intention to deceive [47].
+Deepfake content can spread widely online, and mislead viewers, contributing to the growing issue of information disorder.
+A revealing feature of audio-visual deepfakes is the presence of incongruities between the visual and audio signals, which are, however, increasingly hard to detect due to the quality and realism of current deepfakes.
 This necessitates the development of robust AI-based deepfake detection methods.
-Deepfake detection is a growing area of research [54, 51, 46].
-Existing approaches primarily focus on pixel-level analysis, which examines inconsistencies in gradient variations [65], color variations [30], or artifacts introduced during the generation process [11].
-Also, feature-based methods analyze facial landmarks [72], temporal coherence [27], and other visual cues to identify inconsistencies in deepfake content [40].
-Furthermore, there has been growing interest in detecting and localizing deepfakes in audio-visual content, with both tasks being particularly challenging due to the complex interplay between the two modalities.
-Key approaches include audio-visual synchronization that analyzes the consistency between visual and auditory cues namely face movements and speech [26, 9, 14], feature reconstruction learning with cross-reconstruction similarity estimation to detect anomalies prevalent in fake videos [76], and self-supervised approaches relying on the synchronization patterns learned from real videos only [28, 18].
+Deepfake detection is a growing research area [55].
+Pertinent methods focus on pixel-level (gradient [69], color [31], and artifact analysis [11]) and feature-based (facial landmarks [76], temporal coherence [28]) inconsistencies.
+An expanding area of study involves the detection and localization of deepfakes in audio-visual data, which is especially hard due to the way sound and visuals are intertwined.
+Key approaches analyze the consistency between raw visual and auditory cues [27, 9, 15], consider feature reconstruction learning modules [80], and follow self-supervised approaches learning synchronization patterns from real videos only [29, 19].
+(a) Example of differences between Visual Speech Recognition (VSR) and Audio Speech Recognition (ASR) outputs.
+File path: AV-Deepfake1M/val/id01003/9nmM17wsxGU/00013/fake_video_fake_audio.mp4
+Figure 2: Identifying machine perception discrepancies between visual and audio speech for deepfake detection.
+In (a), a video’s visual and audio streams are separately processed by VSR and ASR models, then the outputs’ normalized Levenshtein distance $\mathtt{d_{L}}$ is calculated.
+In (b,c,d) the $\mathtt{d_{L}}$ distributions are illustrated for FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8].
 Humans integrate multimodal sensory information, such as visual and auditory input, to extract meaningful features and perform a wide range of recognition tasks.
@@
 When visual information conflicts with auditory input, as in the McGurk effect [50] or in the case of deepfakes, increased prediction errors are observed, accompanied by significant changes in brain activity, which often manifest as higher-frequency neural oscillations and localized activation patterns, indicative of heightened cognitive effort [4].
-(a) Example of speech recognition differences between lip-reading and speech-to-text models.
-(b) Aggr. FakeAVCeleb [34].
-(c) Aggr. LAV-DF [10].
-(d) Aggr. AV-Deepfake1M [8].
-Figure 1: The process of identifying audio-visual inconsistency for deepfake detection.
-In (a), a video’s visual and audio components are separately processed by lip-reading and speech-to-text models, then a difference score is calculated.
-In (b,c,d) the difference score distributions are illustrated in the form of violin plots for three different datasets.
-We illustrate how state-of-the-art AI analysis tools could emulate this process in Figure 1.
-In Figure 1(a), an input video is first decomposed into its constituent visual and audio streams.
-These are then fed into state-of-the-art models for lip-reading and speech-recognition [44], and the resulting textual outputs are compared [^1] Python’s $\mathtt{difflib}$ is used (cf.
-https://docs.python.org/3/library/difflib.html)..
-Their difference is computed using the normalized Levenshtein distance, i.e.
-the ratio between the edit distance $\mathtt{delta}$ of the two sentences and the sum of their lengths, which results in scores between 0% (identical) and 100% (completely different).
-Figures 1(b), 1(c) and 1(d) illustrate aggregate results through violin plots computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [34], LAV-DF [10], and AV-Deepfake1M [8].
-On average, lower difference is observed in real videos compared to fake ones across all datasets, indicating the suitability of this score as a measure of incongruity between audio and video.
-Even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which render naive thresholding on this score unsuitable for robust detection.
-This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, while samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
-Motivated by the above, we propose Discourse Modality-information Differentiation (DiMoDif), a deep learning-based deepfake detection architecture that exploits the inter-modality differences in machine perception of speech.
-DiMoDif first decomposes the input video into its visual and audio streams, and extracts the respective video and audio representations based on state-of-the-art lip reading and speech-to-text pre-trained models [44].
-At its core, DiMoDif employs a Transformer encoder architecture [66] with a feature pyramid scheme and local attention configuration, to spot frame-level cross-modal incongruities between the two modality-specific representations, and in that way detect and localize the deepfake forgeries.
-Note that low difference scores frequently occur in partly manipulated fake videos (e.g.
-Figure 1(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level.
-In addition, high difference scores occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf.
+We illustrate how state-of-the-art AI analysis tools could emulate this process in fig.
+2.
+In fig.
+2(a), an input video is decomposed into its visual and audio streams feeding state-of-the-art models for Visual and Audio Speech Recognition (VSR, ASR) [^1] Also known as lip-reading and speech-to-text, respectively.
+[46].
+The difference between the two textual outputs is quantified using the normalized Levenshtein distance ($\mathtt{d_{L}}$), calculated by dividing the edit distance (delta) [^2] Determined through Python’s difflib.
+by the combined length of the texts, resulting in a percentage from 0% (identical) to 100% (completely different).
+Figures 2(b), 2(c) and 2(d) illustrate aggregate results computed on the evaluation sets of three popular audio-visual deepfake detection benchmarks, FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8].
+Real videos, on average, showed lower difference scores $\mathtt{d_{L}}$ than fake videos across all datasets, confirming the score’s ability to measure incongruity.
+However, even though these differences are significant on FakeAVCeleb, they are smaller on LAV-DF, and much smaller on AV-Deepfake1M, which renders naive thresholding unsuitable for robust detection.
+This is due to the fact that FakeAVCeleb’s videos are fully manipulated, while LAV-DF’s samples are manipulated only in parts, and samples from AV-Deepfake1M are manipulated in even smaller parts (on average half the length of those in LAV-DF).
+To leverage machine perception discrepancies between visual and audio speech, we introduce Discourse Modality-information Differentiation (DiMoDif).
+The proposed architecture decomposes the input video into its visual and audio streams, and extracts features via the Discourse-related Feature Extraction (DiFE) module, which uses the pre-trained VSR and ASR models proposed in [46].
+Subsequently, the Modality-information Differentiation (MiD) module, implemented via a Transformer encoder with local cross-modal attention and feature pyramids, identifies frame-level inconsistencies and utilizes them for deepfake detection and localization.
+Note that low difference scores $\mathtt{d_{L}}$ frequently occur in partly manipulated fake videos (e.g.
+fig.
+2(a)); however, DiMoDif manages to classify them as fake by capturing audio-visual divergence at the frame level.
+In addition, high difference scores $\mathtt{d_{L}}$ occasionally occur in real videos; however, DiMoDif leverages alternative features, e.g., hard phonemes [6], to classify them as real (cf.
 section 5.2).
-The architecture is optimized using a composite loss function that accounts for frame-specific detections, overlaps between predicted and ground truth fake intervals, and divergence in corresponding boundaries.
+A composite loss optimizes frame-level detection and fake interval prediction (overlap and boundaries).
 We evaluate DiMoDif on Deepfake detection (DFD) and Temporal Forgery Localization (TFL) tasks using three audio-visual deepfake detection benchmarks, FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8].
-DiMoDif outperforms the state-of-the-art on FakeAVCeleb by +2.8% AUC, exhibits a significant performance increase of +30.5% AUC and +47.88% AP@0.75 on AV-Deepfake1M, and performs on-par on LAV-DF dataset.
-We also provide an analysis of DiMoDif’s generalization abilities, and an extensive ablation analysis to ground our methodological choices.
+DiMoDif outperforms the state-of-the-art by a considerable margin on the challenging AV-Deepfake1M (+30.5% AUC for DFD and +47.88% AP@0.75 for TFL), while it maintains top performance on FakeAVCeleb and LAV-DF.
+The main contributions of this paper are:
+  1. 1.
+    A novel audio-visual deepfake detection methodology leveraging speech representations from both modalities to identify local cross-modal incongruities, called DiMoDif.
+  2. 2.
+    An extensive evaluation analysis of ablation, comparative, robustness, generalization, and in-the-wild experiments demonstrating DiMoDif’s effectiveness on deepfake detection and localization.
+  3. 3.
+    An inherent capability for interpreting DFD outputs through DiMoDif’s cross-modal representation incongruity.
 2 Related Work
+2.1 Deepfake Detection Approaches
+Deepfake detection approaches considering visual-only manipulations [65, 60, 82, 40, 1, 18, 57, 66, 41], while valuable, overlook the audio component of videos and its interplay with the visual content.
+We compare DiMoDif with a representative set of these approaches following previous practices.
+Recent deepfake detection approaches employ multimodal techniques like fusion, contrastive learning, and self-supervised learning to identify discrepancies between visual and audio streams.
+While these methods effectively capture audio-visual inconsistencies, they neglect the speech component of videos.
+For instance, attention-based frameworks analyze synchronization patterns [84], spatio-temporal models assess inter- and intra-modal disharmonies [75], and forgery-aware adaptation is applied to pre-trained ViTs [52].
+Similarly, [58, 32, 86, 72] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms.
+UMMAFormer [80] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives.
+Contrastive learning approaches compare inter-modal representations within a video [15, 14] or between videos [81].
+[27] identifies audio-visual inconsistency by focusing on feature similarity, and [61] by comparing the video’s lip movements with audio-derived lip sequences.
+Self-supervised learning addresses the generalization issue of deepfake detection by exploiting natural audio-visual synchronization in real videos [79, 26, 29, 23, 19, 54, 77, 22].
+Representation learning has also been tailored to modality emotion consistency [48] and articulatory/lip movement consistency [73].
+Finally, works [44, 51, 64, 78, 5] dealing with Temporal Action Localization (TAL) are considered relevant by the deepfake detection literature.
+Related works are presented in [7, 42].
+These rely on a simple threshold on the divergence between VSR and ASR model outputs.
+However, as shown in fig.
+2 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M.
+In contrast, our approach extracts VSR and ASR content representations that are then processed by a Transformer-based fusion mechanism capable of identifying discriminative speech features on top of speech divergence ones.
+Pre-trained VSR models have been utilized as feature extractors by Haliassos et al.
+[30] for visual-only deepfake detection but neither audio speech features nor cross-modal incongruities were considered.
+Finally, [59] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model.
 2.2 Deepfake Detection Datasets
+content manip. target
+dataset size A V A V A V joint task
+DFDC [21] 128,154 DFD
+KoDF [38] 237,942 DFD
+FakeAVCeleb [36] 21,544 DFD
+LAV-DF [10] 136,304 DFD,TFL
+AV-Deepfake1M [8] 1,146,760 DFD,TFL
+Table 1: Deepfake detection datasets providing both video and audio.
+Size determines the number of samples, content specifies the provided modalities, manipulation specifies which modality has been manipulated, target specifies for which modality the dataset provides ground truth, while task is reported in the last column.
 During the past years, several datasets have been proposed for video deepfake detection.
@@
 Finally, AV-Deepfake1M [8] is the largest audio-visual deepfake detection benchmark to date containing over 1.1M real and fake videos with manipulations and targets on both modalities, for both the DFD and TFL tasks.
-In our experimental analysis, we consider FakeAVCeleb, LAV-DF, and AV-Deepfake1M datasets, which meet all requirements of the audio-visual deepfake detection task.
-We also consider DFDC and KoDF for cross-dataset generalization evaluation experiments.
-content manip. target
-dataset size A V A V A V joint task
-DFDC [20] 128,154 DFD
-KoDF [36] 237,942 DFD
-FakeAVCeleb [34] 21,544 DFD
-LAV-DF [10] 136,304 DFD,TFL
-AV-Deepfake1M [8] 1,146,760 DFD,TFL
-Table 1: Deepfake detection datasets providing both video and audio.
-Size determines the number of samples, content specifies the provided modalities, manipulation specifies which modality has been manipulated, target specifies for which modality the dataset provides ground truth, while task is reported in the last column.
-2.2 Deepfake Detection Approaches
-Many deepfake detection approaches consider only video-level manipulations and processing [62, 58, 78, 38, 1, 17, 55].
-Although these are only partly related to DiMoDif  we compare our method with them following previous practice.
-Recent works in deepfake detection have focused on leveraging both visual and audio information through learning mechanisms based on multimodal fusion, contrastive learning, and self-supervised-learning, aiming to uncover cross-modal inconsistencies in deepfake videos.
-Specifically, [80] demonstrated the effectiveness of an attention-based detection framework that identifies the violation of cross-modal synchronization patterns in deepfake videos, while [71] proposed a framework exploiting audio-visual inconsistency by encoding spatio-temporal information, fusing multimodal features, and then classifying based on inter- and intra-modal disharmony levels.
-Similarly, [56, 31, 82] propose joint representation learning frameworks that consider inter- and intra-modal encoding mechanisms.
-UMMAFormer [76] adopts feature reconstruction and cross-reconstruction attention combined with an enhanced feature pyramid network, while BA-TFD & BA-TFD+ [10, 9] adopt 3DCNN and multi-scale vision Transformer backbones guided by contrastive, frame classification, and boundary matching objectives.
-The contrastive learning approach is also adopted by other works for comparing inter-modal representations within a video [14] or between different videos [77].
-Other papers focus on feature similarity to identify audio-visual inconsistencies [26], and [59] compares features of the video’s lip sequence with a lip sequence generated from the audio input.
-Self-supervised approaches that leverage intrinsic synchronization of real videos have also been utilized [75, 25, 28, 22, 18, 53, 73, 21] to tackle the generalization challenge faced by methods trained on specific deepfake generation models.
-A similar motivation with ours is presented in [7, 39], which rely on a simple threshold on the divergence between visual and audio speech recognition model outputs.
-However, as shown in Figure 1 this simple technique is insufficient on benchmarks with harder samples than FakeAVCeleb, such as LAV-DF and AV-Deepfake1M.
-In contrast, our approach extracts lip-reading and speech-to-text content representations that are then processed by a Transformer-based fusion mechanism that is capable of identifying discriminative speech features on top of speech divergence ones.
-Pre-trained lip-reading models have also been utilized as feature extractors in [29], but for visual-only deepfake detection.
-In addition, [57] sets a threshold on cross-modal feature similarity deriving from an audio-visual speech self-supervised representation learning model.
-Also, specialized representation learning has been explored by [47] that considers modality emotion consistency and [69] that considers articulatory and lip movement consistency.
-Finally, several works [42, 50, 61, 74, 5] of the Temporal Action Localization (TAL) domain are considered as relevant by the deepfake detection literature.
 3 Methodology
+In this section we elaborate on the main components of DiMoDif, namely Discourse-related Feature Extraction (DiFE), Modality-information Differentiation (MiD), classification & regression heads, and objective function.
+DiMoDif’s architecture is illustrated in Figure 3.
+Figure 3: The DiMoDif architecture.
 3.1 Problem Formulation
-Consider a dataset $\mathfrak{D}=\{\mathfrak{v}_{i},y_{i},\mathbf{y}_{i}\}_{i=1}^{N}$ where $\mathfrak{v}_{i}=\{v,a\}$ denotes a video, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^3] Utterances mainly comprise monophonic signals, besides conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples.
-$y_{i}\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}_{i}=\{y^{v}_{i,1},\dots,y^{v}_{i,f},y^{a}_{i,1},\dots,y^{a}_{i,f}\}$, with $y^{m}_{i,j}=(d^{m,s}_{i,j},d^{m,e}_{i,j},a^{m}_{i,j})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{i,j}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $j$, along with the corresponding distances $d^{m,s}_{i,j},d^{m,e}_{i,j}\in\mathbb{R}$ between the frame $j$ and the onset $s^{m}_{i,j}$ and offset $e^{m}_{i,j}$ of the fake interval, if the frame $j$ is fake.
-We train deep networks $\mathcal{F}$, the architecture of which is analysed in Section 3.2, on the DFD task $\hat{y}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ and the TFL task $\hat{\mathbf{y}}_{i}=\mathcal{F}(\mathfrak{v}_{i})$ separately, while the predictions $\hat{\mathbf{y}}_{i}$ are decoded by:
-$\hat{a}^{m}_{i,j}>0.5\text{, }\hat{s}^{m}_{i,j}=j-\hat{d}^{m,s}_{i,j}\text{, }\hat{e}^{m}_{i,j}=j-\hat{d}^{m,e}_{i,j}$ (1)
-$\hat{\cdot}$ denotes prediction while its absence ground truth. Henceforth, we omit sample index $i$ to reduce notation complexity.
-3.2 Architecture
-Our architecture is illustrated in Figure 2.
-From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the lip-reading and speech-to-text models proposed in [44].
-Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected.
-For that reason, the proposed architecture that identifies such patterns is named Discourse Modality-information Differentiation (DiMoDif).
-The discourse features, $\mathbf{v}$ and $\mathbf{a}$, are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers.
-Then, we add trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ and employ a separation token $\mathbf{s}\in\mathbb{R}^{d}$ to encode modality information.
-An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{j}\in\mathbb{R}^{d}$, with $j=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [66].
+Consider $(\mathfrak{v},y,\mathbf{y})\in\mathfrak{D}$ where $\mathfrak{v}=\{v,a\}$ denotes a video of dataset $\mathfrak{D}$, containing the visual $v\in\mathbb{R}^{h\times w\times c\times f}$ and the audio $a\in\mathbb{R}^{s}$ signal [^4] Utterances mainly comprise monophonic signals; conversion to single-channel is straightforward if required., with $h$ denoting the height, $w$ the width, $c$ the number of channels, $f$ the number of video frames, and $s$ the number of audio samples.
+$y\in\{0,1\}^{2}$ denotes the Deepfake Detection (DFD) target per modality, while $\mathbf{y}=\{y^{v}_{1},\dots,y^{v}_{f},y^{a}_{1},\dots,y^{a}_{f}\}$, with $y^{m}_{\phi}=(d^{m,s}_{\phi},d^{m,e}_{\phi},a^{m}_{\phi})$, denotes the Temporal Forgery Localization (TFL) target, assigning a forgery ground truth $a^{m}_{\phi}\in\{0,1\}$ to each modality $m\in\{v,a\}$ and frame $\phi$, along with the corresponding distances $d^{m,s}_{\phi},d^{m,e}_{\phi}\in\mathbb{R}$ between frame $\phi$ and the onset $s^{m}_{\phi}$ and offset $e^{m}_{\phi}$ of a fake interval containing $\phi$.
+We train deep networks $\mathcal{F}$ on the DFD task $\hat{y}=\mathcal{F}(\mathfrak{v})$ and the TFL task $\hat{\mathbf{y}}=\mathcal{F}(\mathfrak{v})$ separately, while predictions $\hat{\mathbf{y}}$ are decoded by:
+$\hat{a}^{m}_{\phi}>0.5\text{, }\hat{s}^{m}_{\phi}=\phi-\hat{d}^{m,s}_{\phi}\text{, }\hat{e}^{m}_{\phi}=\phi-\hat{d}^{m,e}_{\phi}$ (1)
+$\hat{\cdot}$ denotes prediction while its absence ground truth.
+3.2 Discourse-related Feature Extraction (DiFE)
+Discourse modality-information differentiation (DiMoDif) is learned by utilizing speech features extracted with disjoint-only encoding of visual and audio streams.
+Otherwise, e.g., by utilizing Audio-Visual Speech Recognition (AVSR) encoders, information missing from one modality due to deepfake artifacts is retrieved from the other resulting in indistinguishable cross-modal differences between real and fake samples.
+From each video $\mathfrak{v}=\{v,a\}$, we extract the visual $\mathbf{v}\in\mathbb{R}^{f\times d_{0}}$ and audio $\mathbf{a}\in\mathbb{R}^{f\times d_{0}}$ features corresponding to its visual $v$ and audio $a$ components, denoting with $d_{0}$ the embedding dimension, based on the distinct VSR and ASR models proposed in [46].
+Then, $\mathbf{v}$ and $\mathbf{a}$ are projected to $d$ dimensions using a lightweight 1-D convolutional network with two ReLU-activated layers.
+Under ideal conditions, in real videos, $\mathbf{v}$ and $\mathbf{a}$ contain the same information corresponding to the discourse, while in deepfakes that contain manipulated parts, differentiation is expected due to artifacts.
+3.3 Modality-information Differentiation (MiD)
+Trainable sequence encoding $S_{v},S_{a}\in\mathbb{R}^{d}$ is added to $\mathbf{v}$ and $\mathbf{a}$, while a separation token $\mathbf{s}\in\mathbb{R}^{d}$ is employed to encode modality information.
+An additional classification token $\mathbf{c}\in\mathbb{R}^{d}$ is only employed in DFD experiments, and position encoding $P_{\tau}\in\mathbb{R}^{d}$, with $\tau=1,\dots,2\cdot f+e$, is applied on all tokens, resulting in tensor $\mathbf{t}=\oplus\{\mathbf{c},\mathbf{v},\mathbf{s},\mathbf{a}\}\in\mathbb{R}^{(2\cdot f+e)\times d}$, with $\oplus$ denoting concatenation, being the input to the Transformer encoder $\mathcal{T}$ [70].
 $e$ depends on the task, being equal to 2 in DFD experiments and equal to 1 in TFL experiments where $\mathbf{c}$ is absent.
-Then, we extract $d$-dimensional feature pyramids corresponding to all $2\cdot f$ visual and audio tokens deriving from all $l$ layers of the Transformer encoder, denoted by $\mathbb{R}^{(2\cdot f)\times l\times d}\ni\mathbf{z}\subset\mathcal{T}(\mathbf{t})$.
+Given that short-term inter-modality inconsistency detection is key to our tasks, multimodal local attention is preferable in contrast to global; besides, it is computationally more efficient.
+Thus, we limit the Transformer’s attention to a multimodal local window of $2\cdot q$ tokens $\tau\in[1+\phi-\left\lfloor q/2\right\rfloor,1+\phi+\left\lfloor q/2\right\rfloor]\cup[f+2+\phi-\left\lfloor q/2\right\rfloor,f+2+\phi+\left\lfloor q/2\right\rfloor]$ for frames $\phi$, while $\mathbf{c}$ and $\mathbf{s}$ attend to all tokens.
+Then, we extract $d$-dimensional feature pyramids corresponding to $\mathbf{c}$ (if applicable), $\mathbf{v}$, and $\mathbf{a}$ tokens deriving from the $l$ layers of the Transformer $\mathcal{T}$, denoted by $\Phi\subset\mathcal{T}(\mathbf{t},q)$, with $\Phi\in\mathbb{R}^{(2\cdot f+e-1)\times l\times d}$.
 To do so, $\mathcal{T}$ processes $\mathbf{t}$ as:
-$\tilde{\mathbf{z}}_{j}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{j-1}))$ (2)
-$\mathbf{z}_{j}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{j}))+\tilde{\mathbf{z}}_{j}$ (3)
-where $j=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [66] with $r$ number of heads, LN denotes Layer Normalization [37], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$.
-Then, $\mathbf{z}=\oplus\{\mathbf{z}_{j}\}_{j=1}^{l}$, omitting the tokens that correspond to $\mathbf{c}$ and $\mathbf{s}$.
-Given that short-term inter-modality inconsistency detection is key to our tasks, local attention is preferable in contrast to global; besides, it is computationally more efficient.
-Thus, we limit the Transformer’s attention to a local window $[j-q/2,j+q/2]$ of $q$ tokens around each token $j$.
-Figure 2: The DiMoDif architecture.
-Finally, in DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\mathbf{z}$, denoted by $\mathbf{z}^{\prime}\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
-$\mathbf{z}^{\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime}\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
-$\mathbf{z}^{\prime\prime\prime}$ $=\text{ReLU}\big{(}\text{LN}(\mathbf{z}^{\prime\prime}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
-$\hat{y}$ $=\mathbf{z}^{\prime\prime\prime}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
+$\tilde{\mathbf{z}}_{\lambda}$ $=\text{MSA}(\text{LN}(\mathbf{z}_{\lambda-1}))$ (2)
+$\mathbf{z}_{\lambda}$ $=\text{MLP}(\text{LN}(\tilde{\mathbf{z}}_{\lambda}))+\tilde{\mathbf{z}}_{\lambda}$ (3)
+where $\lambda=1,\dots,l$ denotes the layer index, $\mathbf{z}_{0}=\mathbf{t}$ is the input, MSA denotes Multi-head Softmax Attention [70] with $r$ number of heads, LN denotes Layer Normalization [39], and MLP denotes a Multi-Layer Perceptron with internal dimensionality $d\cdot u$.
+Then, $\Phi=\oplus\{\mathbf{z}_{\lambda}\}_{\lambda=1}^{l}$, omitting the separation token $\mathbf{s}$.
+3.4 Classification and Regression Heads
+In DFD experiments, predictions $\hat{y}$ are made by a feed-forward classification head $\mathcal{Q}$ that takes as input the classification tokens of $\Phi$, denoted by $\Psi\in\mathbb{R}^{l\times d}$, and consists of three linear layers, the first two of which are followed by Layer Normalization (LN) and ReLU activation:
+$\dot{\Psi}$ $=\text{ReLU}\big{(}\text{LN}(\Psi\cdot\mathrm{W}_{1}+\mathrm{b}_{1})\big{)}$ (4)
+$\ddot{\Psi}$ $=\text{ReLU}\big{(}\text{LN}(\dot{\Psi}\cdot\mathrm{W}_{2}+\mathrm{b}_{2})\big{)}$ (5)
+$\hat{y}$ $=\ddot{\Psi}\cdot\mathrm{W}_{3}+\mathrm{b}_{3}$ (6)
 […]
 3.5 Objective Function
-For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs.
+For the DFD task, we consider the binary cross-entropy loss $\mathfrak{L}_{ce}$ [25], optimizing the real vs.
 fake objective.
-For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals.
-Detailed mathematical expressions explaining the computation of both loss functions can be found in the Appendix A.
+For the TFL task, we propose a composite loss as the combination of three loss functions, namely (1) the focal loss $\mathfrak{L}_{f}$ [45] to optimize the classification objective, while accounting for class imbalance being prevalent in each video’s frames, (2) the DIoU loss $\mathfrak{L}_{d}$ [83] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals, and (3) the smooth L1 loss $\mathfrak{L}_{1}^{s}$ [24] to minimize the distance between predicted and actual boundaries of fake time intervals.
+Detailed mathematical expressions of loss function computation can be found in the supplementary material.
 4 Experimental Setup
@@
 We consider FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8] datasets for training and evaluation.
-Appendix B provides details on class and split sizes.
-FakeAVCeleb contains 21K fake and 500 real samples, thus for balance we added 20K more from VoxCeleb2 dataset [15] including 10K of male and 10K of female speakers of several ethnicities talking in English [^4] Language identification conducted with https://github.com/speechbrain/speechbrain., reaching a total of 41,544 samples randomly split based on a 80-10-10 identity separation.
+FakeAVCeleb contains 21K fake and 500 real samples which we split in 70% for training and 30% for testing, following [75, 54] [^5] To balance classes during training only, we add real samples from VoxCeleb2 [16] including equal number of male and female speakers of several ethnicities talking in English.
+Language identification via https://github.com/speechbrain/speechbrain., while we keep 5% of the training set for validation.
 LAV-DF has been released with standard splits of 78K training, 31K validation, and 26K test samples.
-AV-Deepfake1M has 746K training, 57K validation, and 343K test samples the metadata of which have not been released.
-We also evaluate DiMoDif on the test set of DFDC [20] and a randomly selected 6K-sized sample from KoDF [36] proportionally spanning all synthesis methods and real samples, to assess generalization.
+AV-Deepfake1M has 746K training, 57K validation, and 343K test samples; the metadata of the latter have not been released.
+Supplementary material provides further details on class and split sizes.
 4.2 Evaluation
@@
 We report metrics per dataset in accordance to previous works for comparability.
-On AV-Deepfake1M we evaluate through Codabench [^5] https://deepfakes1m.github.io/evaluation.
-Competing methods are audio-visual deepfake detectors, face manipulation detectors, and temporal action localization models, discussed in Section 2.
+On AV-Deepfake1M we evaluate through Codabench platform [^6] https://deepfakes1m.github.io/evaluation.
+Competing methods are audio-visual deepfake & face manipulation detectors, and temporal action localization models (cf.
+Section 2).
+Also for DFD the naive thresholding on $\mathtt{d_{L}}$ approach (cf.
+fig.
+2).
 4.3 Implementation Details
 We train DiMoDif for 100 epochs with early stopping patience 10 and checkpoint based on validation metrics, AUC for DFD and the sum of AP@$p$ and AR@$n$ for TFL.
-Batch size is set to 64, initial learning rate to 0.001 with ReduceLROnPlateau scheduler [^6] https://pytorch.org/docs/stable/generated/torch.optim.lr_scheduler.ReduceLROnPlateau.html and Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$.
-Also, we set a maximum sequence length $f$ to 600 and zero-pad the smaller.
+Batch size is set to 64, initial learning rate to 0.001 (reduced on plateau) with Adam optimizer, local attention window size $q$ to 15, and focal loss’ hyperparameters $\alpha=0.98$ and $\gamma=2$.
+We set a maximum sequence length $f$ to 600 and zero-pad the smaller.
 We consider a grid for $\mathcal{T}$’s size, with $(d,r,u)\in\{(32,2,1),(64,2,1),(64,4,1),(128,4,2),(256,8,2)\}$ and number of layers $l\in\{1,3,5\}$.
@@
 5 Results
-5.1 Temporal Forgery Localization
-Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
-MDS [14] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
-AGT [50] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
-BMN [42] $\mathcal{V}$ 24.0 7.6 0.1 53.3 41.2 31.6 26.9
-BMN (I3D) [42] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
-AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
-ActionFormer [74] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
-BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
-BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
-UMMAFormer [76] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
-DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
-Table 2: Temporal forgery localization results on LAV-DF [10].
+5.1 Ablation Study
+Other speech-related pre-trained models are equivalent for use as backbones in DiMoDif, e.g., AV-Hubert [63] (cf.
+section 3.2).
+To study their impact, we train DiMoDif on FakeAVCeleb and LAV-DF w/ AV-Hubert features and present the results in Table 2.
+This variant performs slightly worse on FakeAVCeleb and LAV-DF (DFD), better wrt AP@0.5 and worse wrt AR@100 on LAV-DF (TFL).
+Henceforth, we opt for the better performing and more intuitive speech prediction variant (Ma 2022) [46].
+Figure 4 presents ablations on AV-Deepfake1M [^7] We report validation performance as conducting the ablations through Codabench would be impractical..
+Figure 4(a) indicates that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$.
+Also, the feature pyramid scheme provides a small performance increase as indicated by Figure 4(b).
+The size of $\mathcal{T}$ is important for achieving maximum performance as shown in Figure 4(c), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$: the larger the model the higher its performance.
+Supplementary material provides further TFL task ablations on AV-Deepfake1M wrt hyperparameter $\alpha$ and learning rate scheduler, the same analysis on LAV-DF, and a pertinent DFD task ablation analysis.
+FAVC L-DF (DFD) L-DF (TFL)
+AV-Hubert [63] 99.98, 99.35* 99.79, 99.78* 98.00+, 93.00
+(Ma 2022) [46] 99.99, 99.71* 99.94, 99.84* 95.49+, 94.17
+Table 2: Speech-related pre-trained backbone ablation analysis. AP, *AUC, +AP@0.5, AR@100.
+(a) AV-Deepfake1M: Window size $q$
+(b) AV-Deepfake1M: Feature pyramid (FP) use
+(c) AV-Deepfake1M: $\mathcal{T}$ size; $l$ $(d,r,d\cdot u)$
+Figure 4: Ablation and hyperparameter tuning analysis.
+5.2 Deepfake Detection
+Tables 5, 5 and 5, present comparative results in terms of video-level Deepfake Detection (DFD) on FakeAVCeleb [36], LAV-DF [10], and AV-Deepfake1M [8].
+On FakeAVCeleb, DiMoDif outperforms competitive methods achieving +0.8 ACC, and +0.6 AUC.
+On LAV-DF, DiMoDif performs slightly better than UMMAFormer [80], achieving 99.84 AUC.
+On AV-Deepfake1M dataset, DiMoDif outperforms the state-of-the-art by a significant absolute performance increase of +30.5 AUC [^8] Codabench does not provide ACC for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase..
+Table 6 presents cross-manipulation performance on FakeAVCeleb in which DiMoDif achieves 100.0 AVG-FV AP yet 66.4 AP on RVFA.
+Notably, DiMoDif presents perfect detection rate in face-swaps (100.0 AUC).
+Mouth in face-swaps moves occasionally unnaturally yielding inaccurate lip-reading predictions, while audio is unaffected yielding correct speech-to-text predictions.
+DiMoDif translates this inconsistency into a deepfake detection.
+The naive classifier achieves high performance on FakeAVCeleb but moderate performance on LAV-DF and AV-Deepfake1M; however, it outperforms several methods.
+Finally, challenging real samples with $\mathtt{d_{L}}$ over 0.8 (cf.
+section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt naive distance-thresholding.
+Method Modality ACC AUC
+naive $\mathcal{AV}$ 97.7 93.3
+Xception [60] $\mathcal{V}$ 67.9 70.5
+LipForensics [30] $\mathcal{V}$ 80.1 82.4
+FTCN [82] $\mathcal{V}$ 64.9 84.0
+MDS [15] $\mathcal{AV}$ 82.8 86.5
+AVoiD-DF [75] $\mathcal{AV}$ 83.7 89.2
+ART-AVDF [73] $\mathcal{AV}$ 96.4 98.2
+AVFF [54] $\mathcal{AV}$ 98.6 99.1
+DiMoDif (ours) $\mathcal{AV}$ 99.4 99.7
+Table 3: In-dataset perfromance on FakeAVCeleb [36].
 Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio.
 Bold indicates best and underline second to best performance.
-Method Modality AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@50 AR@30 AR@20 AR@10 AR@5
-Meso4 [1] $\mathcal{V}$ 09.86 06.05 02.22 00.59 38.92 38.91 38.81 36.47 26.91
-MesoInception4 [1] $\mathcal{V}$ 08.50 05.16 01.89 00.50 39.27 39.22 39.00 35.78 24.59
-EfficientViT [17] $\mathcal{V}$ 14.71 02.42 00.13 00.01 27.04 26.99 26.43 23.90 20.31
-TriDet+VideoMAEv2 [61, 67] $\mathcal{V}$ 21.67 05.83 00.54 00.06 20.27 20.23 20.12 19.50 18.18
-TriDet+InternVideo [61, 68] $\mathcal{V}$ 29.66 09.02 00.79 00.09 24.08 24.06 23.96 23.50 22.55
-ActionFormer+VideoMAEv2 [74, 67] $\mathcal{V}$ 20.24 05.73 00.57 00.07 19.97 19.93 19.81 19.11 17.80
-ActionFormer+InternVideo [74, 68] $\mathcal{V}$ 36.08 12.01 01.23 00.16 27.11 27.08 27.00 26.60 25.80
-BA-TFD [10] $\mathcal{AV}$ 37.37 6.34 0.19 0.02 45.55 40.37 35.95 30.66 26.82
-BA-TFD+ [9] $\mathcal{AV}$ 44.42 13.64 0.48 0.03 48.86 44.51 40.37 34.67 29.88
-UMMAFormer [76] $\mathcal{AV}$ 51.64 28.07 7.65 1.58 44.07 43.93 43.45 42.09 40.27
-DiMoDif (ours) $\mathcal{AV}$ 86.93 75.95 28.72 5.43 81.57 80.85 80.25 78.84 76.64
-Table 3: Temporal forgery localization results on AV-Deepfake1M [8].
-Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio.
-Bold indicates best and underline second to best performance.
-Tables 2 and 3 present the performance of our method in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8] datasets, respectively.
-On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50,20}, is second to best wrt AR@10, and third wrt AP@{0.5,0.75,0.95}.
-UMMAFormer [76], BA-TFD+ [9] and ActionFormer [74] exhibit similar performance, while the rest of the models have lower localization abilities.
-On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase.
-The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
 Method Modality AUC
+naive $\mathcal{AV}$ 77.8
 F3-Net [57] $\mathcal{V}$ 52.0
@@
 UMMAFormer [80] $\mathcal{AV}$ 99.8
-DiMoDif (ours) $\mathcal{AV}$ 99.6
+DiMoDif (ours) $\mathcal{AV}$ 99.84
 Table 4: Deepfake detection results on LAV-DF [10].
@@
 Method Modality AUC ACC
+naive $\mathcal{AV}$ 62.8* 77.1*
 Video-LLaMA (13B) E5 [79] $\mathcal{AV}$ 50.7 25.1
@@
 MDS [15] $\mathcal{AV}$ 56.6 59.4
-DiMoDif (ours) $\mathcal{AV}$ 96.3 83.3*
+DiMoDif (ours) $\mathcal{AV}$ 96.3 96.3*
 Table 5: Deepfake detection results on AV-Deepfake1M [8].
@@
 Bold indicates best and underline second to best performance.
-*Validation set accuracy is reported as Codabench only provides AUC score.
+*Computed on validation set.
 Category
-Method Modality Pre-training RVFA FVRA-WL FVFA-WL FVFA-FS FVFA-GAN AVG-FV
+Method Modality Pre-training RVFA FVRA-WL FVFA-FS FVFA-GAN FVFA-WL AVG-FV
 AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
+naive $\mathcal{AV}$ 61.2 65.6 99.2 96.7 98.3 95.6 98.8 95.7 99.3 96.8 98.9 96.2
 Unsupervised AVBYOL [26] $\mathcal{AV}$ LRW [17] 50.0 50.0 73.4 61.3 88.7 80.8 60.2 33.8 73.2 61.0 73.9 59.2
@@
 RealForensics [29] $\mathcal{V}$ LRW [17] - - 88.8 93.0 99.3 99.1 99.8 99.8 93.4 96.7 95.3 97.1
-DiMoDif (ours) $\mathcal{AV}$ 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
-Table 6: Deepfake detection results on FakeAVCeleb [34].
+AVFF [54] $\mathcal{AV}$ LRS3 [3] 93.3 92.4 94.8 98.2 100. 100. 99.9 100. 99.4 99.8 98.5 99.5
+DiMoDif (ours) $\mathcal{AV}$ 66.4 51.6 100. 99.8 100. 100. 100. 100. 100. 100. 100. 99.9
+Table 6: Cross-manipulation performance on FakeAVCeleb [36].
 Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio.
@@
 Supervised methods that use pre-training are fine-tuned on FakeAVCeleb, while unsupervised methods are not trained with labels and fake examples.
+5.3 Temporal Forgery Localization
+Method Modality AP@0.5 AP@0.75 AP@0.95 AR@100 AR@50 AR@20 AR@10
+MDS [15] $\mathcal{AV}$ 12.8 1.6 0.0 37.9 36.7 34.4 32.2
+AGT [51] $\mathcal{V}$ 17.9 9.4 0.1 43.2 34.2 24.6 16.7
+BMN (I3D) [44] $\mathcal{V}$ 10.6 1.7 0.0 48.5 44.4 37.1 31.6
+AVFusion [5] $\mathcal{AV}$ 65.4 23.9 0.1 63.0 59.3 54.8 52.1
+ActionFormer [78] $\mathcal{V}$ 95.3 90.2 23.7 88.4 89.6 90.3 90.4
+BA-TFD [10] $\mathcal{AV}$ 76.9 38.5 0.3 66.9 64.1 60.8 58.4
+BA-TFD+ [9] $\mathcal{AV}$ 96.3 85.0 4.4 81.6 80.5 79.4 78.8
+UMMAFormer [80] $\mathcal{AV}$ 98.8 95.5 37.6 92.4 92.5 92.5 92.1
+MMMS-BA [34] $\mathcal{AV}$ 97.5 95.2 39.0 94.0 93.4 95.9 89.4
+DiMoDif (ours) $\mathcal{AV}$ 95.5 87.9 20.6 94.2 93.7 92.7 91.4
+Table 7: Temporal forgery localization results on LAV-DF [10].
+Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio.
+Bold indicates best and underline second to best performance.
+Method Modality AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@50 AR@30 AR@20 AR@10 AR@5
+PyAnnote (Zero-Shot) [56] $\mathcal{A}$ 00.03 00.00 00.00 00.00 00.67 00.67 00.67 00.67 00.67
+Meso4 [1] $\mathcal{V}$ 09.86 06.05 02.22 00.59 38.92 38.91 38.81 36.47 26.91
+MesoInception4 [1] $\mathcal{V}$ 08.50 05.16 01.89 00.50 39.27 39.22 39.00 35.78 24.59
+EfficientViT [18] $\mathcal{V}$ 14.71 02.42 00.13 00.01 27.04 26.99 26.43 23.90 20.31
+TriDet+VideoMAEv2 [64, 71] $\mathcal{V}$ 21.67 05.83 00.54 00.06 20.27 20.23 20.12 19.50 18.18
+ActionFormer+VideoMAEv2 [78, 71] $\mathcal{V}$ 20.24 05.73 00.57 00.07 19.97 19.93 19.81 19.11 17.80
+BA-TFD [10] $\mathcal{AV}$ 37.37 6.34 0.19 0.02 45.55 40.37 35.95 30.66 26.82
+BA-TFD+ [9] $\mathcal{AV}$ 44.42 13.64 0.48 0.03 48.86 44.51 40.37 34.67 29.88
+UMMAFormer [80] $\mathcal{AV}$ 51.64 28.07 7.65 1.58 44.07 43.93 43.45 42.09 40.27
+MMMS-BA [34] $\mathcal{AV}$ 62.75* 35.87* - 18.37* 57.49* - 55.94* 54.28* -
+DiMoDif (ours) $\mathcal{AV}$ 86.93 75.95 28.72 5.43 81.57 80.85 80.25 78.84 76.64
+Table 8: Temporal forgery localization results on AV-Deepfake1M [8].
+Modality denotes the model’s input type with $\mathcal{V}$ being visual and $\mathcal{A}$ audio.
+Bold indicates best and underline second to best performance.
+*Computed on validation set.
+Tables 7 and 8 present the performance of DiMoDif in comparison to state-of-the-art, in terms of Temporal Forgery Localization (TFL), on LAV-DF [10] and AV-Deepfake1M [8], respectively.
+On LAV-DF, DiMoDif outperforms all competitive approaches wrt AR@{100,50}, is second to best wrt AR@{20,10}, and performs below state-of-the-art wrt AP@{0.5,0.75,0.95}.
+MMMS-BA [34], UMMAFormer [80], BA-TFD+ [9] and ActionFormer [78] exhibit similar performance, while the remaining models have lower localization ability.
+On AV-Deepfake1M, DiMoDif outperforms all competitive approaches wrt all metrics exhibiting significant performance increase.
+The proposed method’s AR is almost double compared with the second to best, while it exhibits a notable absolute increase of +35.29% and +47.88% in AP at 0.5 and 0.75, respectively.
+5.4 Interpretability
+Figure 5: Fake video sample (manipulated in red). Avg. and std. (across layers $\lambda$) of cross-modal similarity and frame-level fake probability shown in blue and purple, respectively.
+DiMoDif is inherently interpretable as its frame-level cross-modal representations’ similarity reflects the level of congruity between visual and audio speech.
+Figure 5 illustrates a fake video example from LAV-DF [^9] LAV-DF/test/000010.mp4 with one small manipulated part during which cross-modal similarity is significantly reduced indicating audio-visual incongruity while during real parts similarity is close to 1.0.
+Fake frames are perfectly identified by DiMoDif with $\max_{m}(\hat{a}^{m}_{\phi})>>0.5$.
+More real and fake examples are provided in suppl.
+material.
+5.5 Robustness
+Figure 6: Robustness to visual (Gaussian blur) and audio (Gaussian noise) distortions.
+Figure 6 illustrates DiMoDif’s robustness to visual (Gaussian blurring) and audio (Gaussian noise) distortions at 5 levels of intensity.
+The analysis is conducted on FakeAVCeleb test set, following the evaluation protocol of [54, 23, 30] [^10] Perturbations’ implementation directly sourced from https://github.com/EndlessSora/DeeperForensics-1.0..
+DiMoDif is not affected by the audio perturbations, outperforming AVFF [^11] Code unavailable.
+Value approximations from [54]’s supplementary material.
+[54] by approx.
++ 6.5 AUC at intensity level 5.
+Robustness to visual perturbations is gradually reduced to 96.5 AUC, approx.
++25.0 AUC compared to AVFF.
+Supplementary material shows that DiMoDif is also more robust than AVFF under saturation, contrast, block-wise, and JPEG compression visual perturbations, yet less robust under Gaussian noise and video compression.
+Also, DiMoDif is more robust than AVFF under pitch shift, reverberence, and audio compression audio perturbations, achieving an unprecedented 99.3 AUC on average at intensity level 5 indicating the high robustness of speech prediction features on such distortions.
+5.6 Generalization
+This paper is the first to conduct generalization experiments on the considered datasets.
+Table 12 presents DiMoDif’s cross-dataset performance on DFD.
+The reported performance, although high, indicates that the training data affect generalization due to scale, fake part duration, and generator quality.
+A small pool of low quality easier fake data (FakeAVCeleb) makes the model prone to overfitting reducing its generalization ability to partially manipulated data generated through more complex approaches (LAV-DF, AV-Deepfake1M).
+In contrast, large pools of high-quality and harder deepfakes (LAV-DF, AV-Deepfake1M) result in richer and more robust representations, and less overfitting.
+Supplementary material provides further generalization evaluations.
 Test dataset
-FakeAVCeleb [34] LAV-DF [10] AV-Deepfake1M [8] DFDC [20] KoDF [36]
-ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC ACC AP AUC
-Training dataset FakeAVCeleb [34] 98.1 99.9 99.9 39.3 59.5 61.9 25.3 55.1 55.1 51.4 62.7 60.8 65.5 77.4 32.7
-LAV-DF [10] 43.6 70.7 71.3 95.6 99.6 99.6 32.1 69.7 66.2 51.8 52.9 54.5 43.8 88.8 66.5
-AV-Deepfake1M [8] 56.3 82.0 85.4 59.2 83.2 81.5 87.7 97.3 97.3 51.3 50.5 49.7 77.0 77.7 33.4
-Table 7: Generalization performance on Deepfake detection task.
+FakeAVCeleb LAV-DF AVD1M*
+AP AUC AP AUC AP AUC
+Training dataset FakeAVCeleb 99.99 99.71 93.10 84.47 77.91 54.00
+LAV-DF 99.69 90.25 99.94 99.84 88.46 70.40
+AV-Deepfake1M 99.69 90.69 94.98 86.30 99.72 99.18
+Table 9: Generalization on DFD. * Validation set.
+5.7 Real-world Analysis
+We also present an analysis of DiMoDif’s performance in the wild.
+Initially, we compiled a collection of 204 Internet videos, across several languages (incl.
+Out-Of-Distribution $\mathtt{OOD}$) and varying lengths ($\mathtt{max}$ 46, $\mathtt{median}$ 1), with 75 of them being real and 129 of them being fake.
+Fakes were selected for being popular deepfakes, supporting conspiracy theories, or spreading political misinformation, and real videos for containing discourse (e.g., news anchoring).
+Next, we developed a pre-processing pipeline to address model’s limitations.
+Chunking is considered for accepting any-length videos, while frame groups with small faces (e.g.
+depicting crowds or wide shots), no faces, or small duration (¡2s) are discarded.
+We utilize DiMoDif trained on AV-Deepfake1M (TFL), and evaluate the ratio of fake to valid video parts, $\mathtt{c}=\mathtt{fake}/\mathtt{valid}$, as the model occasionally flags small parts of authentic videos as deepfakes.
+DiMoDif achieves 82.1 AP, indicating robust detection of real-world cases, with linear processing complexity $\mathcal{O}(\mathtt{[sec]})$.
+Figure 7 illustrates the distribution of $\mathtt{c}$ for real vs.
+fake and decision examples.
+For reference, RealForensics [29] achieves 76.4 AP on the same set.
+Additionally, although the developed pipeline is tailored to real-world cases it excels on evaluation benchmarks, i.e., achieves 99.1 AP on FakeAVCeleb, 93.5 AP on LAV-DF*, and 90.6 AP on AV-Deepfake1M* (*1K random unseen samples).
+Further details are provided in supplementary material.
+Figure 7: Distribution of coverage $\mathtt{c}$ for in-the-wild real vs. fake and examples of decisions. Red parts are flagged as fake.
+6 Conclusions
+In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech.
+It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities.
+It considers a visual and audio speech recognition feature extraction stage, and a hierarchical cross-modal fusion learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries.
+An extensive evaluation study indicates the effectiveness of our framework exhibiting state-of-the-art performance.
+Limitations: DiMoDif’s language coverage is constrained by its speech recognition backbones although it was found to generalize well to Korean (cf.
+supplementary material) and in-the-wild $\mathtt{OOD}$ examples, while performance may be degraded by modality misalignment in real-world samples.
+Additionally, the trained model assumes a single, consistently visible speaker, sufficient video quality for facial landmark detection, and the availability of both visual and audio modalities.
+Development for the real-world analysis has addressed in part some of these limitations through video pre-processing.
+We aim to further improve in-the-wild applicability and extend functionality to multi-speaker scenarios.
+7 Acknowledgments
+This work has been partially funded by the Horizon Europe projects AI4Trust (GA No. 101070190) and vera.ai (GA No. 101070093).
+Appendix A Objective function details
+For the DFD task, we consider the binary cross-entropy loss [25], directly optimizing the real vs. fake objective:
+$\mathfrak{L}_{DFD}=\mathfrak{L}_{ce}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{\lambda=1}^{l}y[m]\text{log}\hat{y}[\lambda,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[\lambda,m])$ (7)
+where $\lambda$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
+For the TFL task, we consider a combination of three loss functions.
+For clarity we omit to also include the symbols for the averaging across $l$ layers, which however takes place as well.
+Specifically, we use:
+  1. 1.
+    The focal loss [45] to optimize the classification objective, while accounting for class imbalance being prevalent among each video’s frames:
+    $\mathfrak{L}_{f}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}-\alpha_{t}^{m,\phi}(1-p_{t}^{m,\phi})^{\gamma}\text{log}(p_{t}^{m,\phi})$ (8)
+    in which:
+    $\alpha_{t}^{m,\phi}=\begin{cases}\alpha&\text{if }y^{m}_{\phi}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
+    $p_{t}^{m,\phi}=\begin{cases}\hat{y}^{m}_{\phi}&\text{if }y^{m}_{\phi}=1\\ 1-\hat{y}^{m}_{\phi}&\text{otherwise}\end{cases}$ (10)
+    while $\alpha$ and $\gamma$ are hyperparameters.
+  2. 2.
+    The DIoU loss [83] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals:
+    $\mathfrak{L}_{d}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{\phi},\mathbf{b}^{m}_{\phi})+\\ \frac{\rho^{2}(\hat{b}^{m}_{\phi},b^{m}_{\phi})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,\phi}$ (11)
+    where $\phi$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{\phi}=[s^{m}_{\phi},e^{m}_{\phi}]$ denotes a time interval, $b^{m}_{\phi}=0.5\cdot(e^{m}_{\phi}+s^{m}_{\phi})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,\phi}$ is an indicator function that denotes if frame $\phi$ of modality $m$ is fake.
+  3. 3.
+    The smooth L1 loss [24] to minimize the distance between predicted and actual boundaries of fake time intervals:
+    $\mathfrak{L}_{1}^{s}=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{\phi}-\hat{s}^{m}_{\phi})+\\ h(e^{m}_{\phi}-\hat{e}^{m}_{\phi})\bigg{)}\cdot\mathds{1}_{m,\phi}$ (12)
+    in which $\mathds{1}_{m,\phi}$ is an indicator function that denotes if frame $\phi$ of modality $m$ is fake, and $h$ is defined as:
+    $h(x)=\begin{cases}0.5x^{2}&\text{if }\lvert x\rvert<1\\ \lvert x\rvert-0.5&\text{otherwise}\end{cases}$ (13)
+We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{\phi=1}^{f}\mathds{1}_{m,\phi}$:
+$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{f}+\mathfrak{L}_{d}+\mathfrak{L}_{1}^{s}\big{)}/p$ (14)
+We finally average both $\mathfrak{L}_{DFD}$ and $\mathfrak{L}_{TFL}$ across batch samples during training.
+Appendix B Dataset class and split sizes
+Table 10 provides details on dataset class and split sizes.
+Although, FakeAVCeleb paper [36] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples (incl.
+500 RVFA & 500 RVRA) instead of the commonly reported 20,000 size for this dataset.
+Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”.
+Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
+Appendix C Further Ablations
+Figure 8 presents the ablations on LAV-DF (TFL) corresponding to the ablations presented in the main manuscript on AV-Deepfake1M, namely for window size $q$, feature pyramid use, and Transformer $\mathcal{T}$ size.
+The conclusions drawn from these ablations on both datasets agree.
+Figure 9 then presents further TFL task ablations on both LAV-DF and AV-Deepfake1M.
+Specifically, Figures 9(a) and 9(b) indicate little to no sensitivity to the focal loss’ hyperparameter $\alpha$.
+Figures 9(c) and 9(d) indicate that the use of a learning rate scheduler is beneficial however the type of scheduler is not significantly affecting the results.
+Next, Figure 10 presents ablations on FakeAVCeleb, LAV-DF, and AV-Deepfake1M wrt the DFD task.
+The results indicate that on DFD a window size of $q=15$ is optimal, but performance is not increased with feature pyramids nor with different learning rate schedules.
+Also, medium-sized models achieve maximum performance.
+Appendix D Interpretability
+Figures 11 and 12 present further interpretability plots for 8 real and 8 fake video samples from LAV-DF.
+Cosine similarity at the frame-level reveals DiMoDif’s decision reasoning in these cases.
+However, there are some rare cases in which DiMoDif correctly identified the modified parts although cross-modal similarity is high, such as in the example presented in Figure 13.
+In such cases DiMoDif is forced to successfully identify alternative manipulation features sacrificing interpretability.
+Appendix E Robustness
+Robustness to visual distortions: Figure 14 illustrates the performance of DiMoDif under unseen visual perturbations (our model has been trained without any visual augmentations) commonly present in real-world scenarios.
+We follow the evaluation protocol of [54, 23, 30] which consider the FakeAVCeleb test set and directly source the implementations of visual perturbations from https://github.com/EndlessSora/DeeperForensics-1.0.
+Specifically, videos are subject to saturation, contrast, block-wise, Gaussian noise, Gaussian blur, JPEG compression, and video compression, while we also compute the average robustness across perturbations.
+DiMoDif clearly outperforms RealForensics [29] under all perturbation scenarios except video compression at intensity levels 4 and 5.
+Also, our method outperforms AVFF [54] under saturation, contrast, block-wise, Gaussian blur, and JPEG compression at all levels of intensity, and is outperformed by AVFF under Gaussian noise and video compression at all levels of intensity.
+Wrt average AUC DiMoDif slightly outperforms AVFF and is much more robust than RealForensics.
+Notably, DiMoDif exhibits exceptional robustness wrt AP on average achieving 99.7 at intensity level 5, indicating near-perfect detection of positive samples despite significant visual distortions.
+In approximately 0.06% of cases, high-intensity Gaussian noise (level 5) prevented landmark detection, making lip-reading impossible and thus precluding the application of our method.
+Robustness to audio distortions: Figure 15 illustrates the performance of DiMoDif under unseen audio perturbations (our model has been trained without any audio augmentations) commonly present in real-world scenarios.
+We follow the evaluation protocol of [54] that considers the FakeAVCeleb test set and implement the audio perturbations using torchaudio, pysndfx, and pydub Python libraries as indicated therein.
+Specifically, the videos are subject to Gaussian noise, pitch shift, reverberence, and audio compression at 5 levels of intensity.
+DiMoDif achieves unprecedented levels of robustness to audio perturbations with 99.3 average AUC across all perturbations at intensity level 5.
+This fact indicates that audio speech recognition is not affected by such perturbations and the corresponding features are effective under any perturbation level scenario.
+In addition, DiMoDif outperforms AVFF under all perturbation scenarios and at all intensity levels.
+[…]
+Appendix F Generalization
+We evaluate DiMoDif (trained on FakeAVCeleb, LAV-DF, and AV-Deepfake1M) on the test set of DFDC [21] and a randomly selected 6K-sized sample from KoDF [38] proportionally spanning all synthesis methods and real samples.
+Table 12 presents the corresponding results.
+Training on FakeAV-Celeb [36] leads to poor performance on DFDC but to high performance on KoDF with 98.8 AP.
+Training on LAV-DF [10] leads to almost random performance on DFDC but high performance on KoDF with 97.9 AP.
+Training on AV-Deepfake1M [8] dataset leads to random performance on DFDC but high performance on KoDF with 98.9 AP.
+In general, very high generalization performance is observed on KoDF, although Korean are not recognized by the (Ma 2022) backbones, indicating the learning of cross-language features by DiMoDif.
+Also, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.2.
+Our models focus on cross-modal speech-related information differences, while many DFDC videos contain multiple talking persons [^12] E.g., DFDC/test/aalyqplqns.mp4, moving persons [^13] E.g., DFDC/test/chyfcacfjr.mp4 or wide shots [^14] E.g., DFDC/test/ljnjhvpnhq.mp4" rendering lip-reading hard, do not contain talking [^15] E.g., DFDC/test/aarpyivfys.mp4 or the talking person is different from the depicted one [^16] E.g., DFDC/test/mfjljhqhca.mp4; thus, visual-only face manipulation detectors lead to better generalization in this dataset.
+Table 13 presents cross-dataset performance on TFL wrt all AP@{0.5, 0.75, 0.90, 0.95} and AR@{100, 50, 30, 20, 10, 5} metrics.
+Cross-dataset performance on TFL is shown to be much more challenging than on DFD.
+Also, the results are even worse when training on LAV-DF in contrast to AV-Deepfake1M which contains half-sized fake parts, thus presenting a more challenging objective, while also being much larger.
+Table 14 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M reaching 99.2 AP & 95.5 AUC (AVG-FV) and 93.9 AP & 95.5 AUC (RVFA).
+Training on LAV-DF also results in good generalization reaching 98.5 AP & 91.2 AUC (AVG-FV) and 81.3 AP & 76.8 AUC (RVFA).
+Note, that this is not a cross-manipulation analysis as the models trained on LAV-DF and AV-Deepfake1M datasets have seen all manipulation types (RVRA, RVFA, FVFA, FVRA).
+Thus, metrics of the 2nd and 3rd rows are comparable but the 1st row (showing cross-manipulation performance) is only provided for reference and is not comparable with the other two.
+Appendix G Real-world analysis
+LABEL:tab:real_world_info provides further information on the set of videos used for the real-world analysis.
+Specifically, we provide the public URL (if exists), the spoken language, the duration of the video, the processing (by DiMoDif) time, and the Coverage score $\mathtt{c}$.
+Language has been automatically retrieved through https://speechbrain.github.io/.
+Finally, 16 of the videos do not have public URLs.
+Appendix H Error analysis
+Figure 16 presents two examples of DiMoDif’s misclassifications.
+Specifically, in the real video shown in fig.
+16(a), a non-verbal vocalization (panting imitation) between 6.3s and 7.3s is incorrectly interpreted as inconsistency, leading DiMoDif to assign a high fake probability to those frames.
+In contrast, fig.
+16(b) features a visual-only manipulated segment from 0.3s to 0.78s; here, despite a slight reduction in cross-modal inconsistency, the high fidelity of the utterance deceives the model.
+FakeAVCeleb LAV-DF AV-Deepfake1M
+train val test total train val test total train val test total
+FVFA 6,996 561 3,278 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
+FVRA 6,363 470 2,876 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
+RVFA 326 24 150 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
+RVRA 13,689 22 160 13,871 21,254 8,271 6,906 36,431 186,666 14,235 - -
+total 27,374 1,077 6,464 34,915 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
+Table 10: Dataset class and split sizes. F: fake, R: real, V: video, A: audio.
+(a) LAV-DF: Window size $q$
+(b) LAV-DF: Feature Pyramid (FP) use
+(c) LAV-DF: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
+Figure 8: Ablation and hyperparameter tuning analysis on LAV-DF (TFL).
+(a) LAV-DF: Focal loss hyperparmeter $\alpha$
+(b) AV-Deepfake1M: Focal loss hyperparmeter $\alpha$
+(c) LAV-DF: Learning rate scheduler
+(d) AV-Deepfake1M: Learning rate scheduler
+Figure 9: Ablation and hyperparameter tuning analysis on LAV-DF and AV-Deepfake1M (TFL).
+(a) FakeAVCeleb: Window size $q$
+(b) LAV-DF: Window size $q$
+(c) AV-Deepfake1M: Window size $q$
+(d) FakeAVCeleb: Feature Pyramid (FP) use
+(e) LAV-DF: Feature Pyramid (FP) use
+(f) AV-Deepfake1M: Feature Pyramid (FP) use
+(g) FakeAVCeleb: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
+(h) LAV-DF: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
+(i) AV-Deepfake1M: $\mathcal{T}$’s size; $l$ $(d,r,d\cdot u)$
+(j) FakeAVCeleb: Learning rate scheduler
+(k) LAV-DF: Learning rate scheduler
+(l) AV-Deepfake1M: Learning rate scheduler
+Figure 10: Ablation and hyperparameter tuning analysis on FakeAVCeleb, LAV-DF, and AV-Deepfake1M (DFD).
+(a) Real #1: LAV-DF/test/000000.mp4
+(b) Real #2: LAV-DF/test/000001.mp4
+(c) Real #3: LAV-DF/test/000005.mp4
+(d) Real #4: LAV-DF/test/000012.mp4
+(e) Real #5: LAV-DF/test/000016.mp4
+(f) Real #6: LAV-DF/test/000032.mp4
+(g) Real #7: LAV-DF/test/000036.mp4
+(h) Real #8: LAV-DF/test/000040.mp4
+Figure 11: Cross-modal cosine similarity on DiMoDif’s representations of 8 real examples.
+(a) Fake #1: LAV-DF/test/000003.mp4
+(b) Fake #2: LAV-DF/test/000008.mp4
+(c) Fake #3: LAV-DF/test/000011.mp4
+(d) Fake #4: LAV-DF/test/000014.mp4
+(e) Fake #5: LAV-DF/test/000015.mp4
+(f) Fake #6: LAV-DF/test/000018.mp4
+(g) Fake #7: LAV-DF/test/000019.mp4
+(h) Fake #8: LAV-DF/test/000030.mp4
+Figure 12: Cross-modal cosine similarity on DiMoDif’s representations of 8 fake examples.
+Figure 13: Non-interpretable fake example (LAV-DF/test/000017.mp4).
+Figure 14: Robustness to unseen visual perturbations. AVFF and RealForensics performance taken from [54]’s supplementary material.
+Figure 15: Robustness to unseen audio perturbations. AVFF performance taken from [54]’s supplementary material.
+Table 11: Text prediction robustness of DiMoDif’s feature extraction backbones (Ma 2022) [46], to visual and audio distortions for the real video sample: FakeAVCeleb/RealVideo-RealAudio/Caucasian (European)/men/id00055/00120.mp4.
+First row presents predictions on the original video while the next ones use versions of it with visual or audio distortions as input.
+$\mathcal{V}$: transcript from visual input, $\mathcal{A}$: transcript from audio input.
+Dist. Lev. Transcript
+- - $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
+$\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+Visual distortions
+Saturation 1 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
+2 $\mathcal{V}$: how i played with growing up and always had great ability great talent from a young age
+3 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
+4 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+5 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+Contrast 1 $\mathcal{V}$: yeah i played with growing up and always had a great ability great talent from a young age
+2 $\mathcal{V}$: how i played with growing up and always had great ability great talent from a young age
+3 $\mathcal{V}$: played with growing up and always had great ability great talent from a young age
+4 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
+5 $\mathcal{V}$: yeah i played with harold growing up and he always had great ability great talent from a young age
+Block-wise 1 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+2 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+3 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+4 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+5 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+Gauss. noise 1 $\mathcal{V}$: i played with carol growing up and he always had a great ability great talent from a young age
+2 $\mathcal{V}$: i played with her growing up and it was a great ability great talent from a young age
+3 $\mathcal{V}$: everybody was scared of growing up with this great ability great talent from a young age
+4 $\mathcal{V}$: our plan was to start growing up and you can see how i beat the grass down from the high
+5 $\mathcal{V}$: -
+Gauss. blur 1 $\mathcal{V}$: i played with growing up and always had great ability great talent from a young age
+2 $\mathcal{V}$: how i played with carol growing up and he always had a great ability great talent from a young age
+3 $\mathcal{V}$: how i played with carol growing up and he always has a great ability great talent from a young age
+4 $\mathcal{V}$: how i played with carol growing up he always has great ability great talent from a young age
+5 $\mathcal{V}$: now i played with carol growing up he always had great ability great talent from a young age
+JPEG compr. 1 $\mathcal{V}$: you know i played with growing up and always had great ability great talent from a young age
+2 $\mathcal{V}$: i played with growing up and always had great ability great talent from a young age
+3 $\mathcal{V}$: now i’m here with the care of growing up and i always like to see the great ability to raise hands for without hands
+4 $\mathcal{V}$: i will play with you the character i’ll be happy once i turn to you i’ll play you as i grew up in high
+5 $\mathcal{V}$: there were pictures of his car flipped up so he said to the red bricks at the right angle that’s
+Video compr. 1 $\mathcal{V}$: yeah i played with growing up and always had great ability great talent from a young age
+2 $\mathcal{V}$: how i played with how i’ve grown up and always had the right ability the right talent from a young age
+3 $\mathcal{V}$: a little bit we’ve now grown up and we’re starting to create building a great town from a young age
+4 $\mathcal{V}$: now i’ll be able to fire from the camera you know adjust it to the front but the right hand will be high
+5 $\mathcal{V}$: i’ll make sure it’s just a narrow volume count it will take you to the right place and the right hand will be high
+Audio distortions
+Gauss. noise 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+Pitch shift 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+2 $\mathcal{A}$: yeah i played with growing up and he always had great ability great talent from a young age
+3 $\mathcal{A}$: you know papers cars growing up and it was a great time for the young age
+4 $\mathcal{A}$: yeah played of gaff growing up and he was a great and it’s a great kind of for a young age
+5 $\mathcal{A}$: -
+Reverberence 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+Audio compr. 1 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+2 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+3 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+4 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+5 $\mathcal{A}$: know i played with growing up and he always had great ability great talent from a young age
+Test dataset
+DFDC [21] KoDF [38]
+AP AUC AP AUC
+Training dataset FakeAVCeleb [36] 60.8 58.9 98.8 69.9
+LAV-DF [10] 52.9 54.5 97.9 56.7
+AV-Deepfake1M [8] 50.5 49.8 98.9 72.8
+Table 12: Generalization on Deepfake detection task.
 AP@0.5 AP@0.75 AP@0.9 AP@0.95 AR@100 AR@50 AR@30 AR@20 AR@10 AR@5
@@
 Table 13: Generalization performance on Temporal Forgery Localization task.
-5.2 Deepfake Detection
-Tables 5, 5 and 6, present comparative results in terms of video-level Deepfake Detection (DFD) on LAV-DF [10], AV-Deepfake1M [8], and FakeAVCeleb [34], respectively.
-On LAV-DF, DiMoDif performs on-par with the state-of-the-art method UMMAFormer [76], achieving 99.6 AUC, better than the remaining competitive methods.
-On AV-Deepfake1M dataset, DiMoDif outperforms all state-of-the-art competitive methods by a significant absolute performance increase of +30.5% in terms of AUC [^7] Accuracy (ACC) is also reported by other papers; however, Codabench does not provide it for the test samples, thus we report the validation accuracy, which, even though is not comparable, still provides a sense on the level of uncalibrated threshold performance increase..
-On FakeAVCeleb, DiMoDif outperforms all competitive methods on average achieving +4.1% absolute AP increase, and +2.8% absolute AUC increase, as well as in most forgery types.
-Finally, challenging real samples with normalized Levenshtein distance over 0.8 (cf.
-Section 1), are scored by DiMoDif with an average of 0.19, and 0.16 on LAV-DF, and AV-Deepfake1M, respectively, attesting to DiMoDif’s advantage wrt a naive distance-thresholding approach.
-5.3 Generalization
-This paper is the first to conduct generalization experiments on the used datasets.
-Table 7 presents DiMoDif’s cross-dataset performance on DFD.
-Training on FakeAVCeleb [34] leads to poor performance on LAV-DF, AV-Deepfake1M, and DFDC but to moderate performance on KoDF with 77.4 AP.
-Training on LAV-DF [10] leads to moderate performance on FakeAVCeleb and AV-Deepfake1M, low performance on DFDC but good performance on KoDF with 88.8 AP.
-Training on AV-Deepfake1M [8] dataset leads to high scores on FakeAVCeleb and LAV-DF with 85.4 AUC and 81.5 AUC respectively, moderate performance on KoDF with 77.7 AP, but low performance on DFDC.
-In general, low generalization is observed on DFDC, which can be partly explained by the discussion in Section 2.1.
-Our models focus on cross-modal information differences, while many DFDC videos do not contain talking or the talking person is different from the depicted one; thus, a visual-only face manipulation detector would likely lead to better results.
-Appendix C provides detailed generalization evaluation on FakeAVCeleb with scores per forgery type.
-Table 8 presents cross-dataset performance on TFL.
-Training on AV-Deepfake1M leads to acceptable performance on LAV-DF, in contrast to training on LAV-DF and evaluating on AV-Deepfake1M.
-5.4 Ablation Analysis
-(a) Window size $q$ (LAV-DF)
-(b) Window size $q$ (AV-Deepfake1M)
-(c) Feature pyramid use (LAV-DF)
-(d) Feature pyramid use (AV-Deepfake1M)
-(e) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
-(f) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
-Figure 3: Ablation and hyperparameter tuning analysis.
-Figure 3 presents an ablation analysis that supports our methodological choices.
-For AV-Deepfake1M we report validation performance as conducting the ablations through Codabench would be impractical.
-These results pertain to TFL, while further ablations as well as the same analysis wrt DFD can be found in Appendix D.
-Figures 3(a) and 3(b) indicate that a small window size $q=15$ for Transformer’s $\mathcal{T}$ local attention is optimal, especially in contrast to attending on all tokens determined by $q=0$.
-Also, the feature pyramid scheme provides a small performance increase as indicated by Figures 3(c) and 3(d).
-The size of $\mathcal{T}$ plays crucial role in achieving maximum performance as shown in Figures 3(e) and 3(f), both in terms of number of layers $l$ and layer size determined by input dimension $d$, number of attention heads $r$, and internal dimension $d\cdot u$.
-Specifically, the larger the model the higher its performance.
-6 Conclusions
-In this work, we propose an audio-visual deepfake detection and localization framework that leverages cross-modal differences in machine perception of speech.
-It hinges on the assumption that the visual and audio signals of videos with real discourse coincide wrt information, in contrast to deepfakes that exhibit cross-modal incongruities.
-It considers a video and audio speech recognition feature extraction stage, and a Transformer encoder-based learning stage along with a combination of three loss functions that optimize frame-level detection, predicted interval overlap, and boundaries.
-An extensive evaluation study indicates the effectiveness of our framework reaching state-of-the-art performance and in several cases surpassing it.
-Our model is limited by the languages recognized by the used speech-recognition backbones, although it exhibits good generalization to Korean.
-Also, it assumes one continually visible speaking person, at least moderate video quality and face size to detect face landmarks, and requires both modalities for manipulation detection.
-Future work includes joint training on Deepfake Detection and Temporal Forgery Localization tasks to assess potential benefits and extension to multiple speakers.
-7 Acknowledgments
-This work has been partially funded by the Horizon Europe projects AI4Trust (GA No. 101070190) and vera.ai (GA No. 101070093).
-Appendix A Objective function details
-For the DFD task, we consider the binary cross-entropy loss [24], directly optimizing the real vs. fake objective:
-$\mathfrak{L}_{DFD}=-\frac{1}{2l}\sum_{m\in\{v,a\}}\sum_{j=1}^{l}y[m]\text{log}\hat{y}[j,m]+\\ (1-y[m])\cdot\text{log}(1-\hat{y}[j,m])$ (7)
-where $j$ is layer index, $m$ is modality index, and $x[\cdot]$ is used to denote vector or matrix element access operation (array indexing).
-For the TFL task, we consider a combination of three loss functions.
-For clarity we omit to also include the symbols for the averaging across $l$ layers, which however takes place as well.
-Specifically, we use:
-  1. 1.
-    The focal loss [43] to optimize the classification objective, while accounting for class imbalance being prevalent among each video’s frames:
-    $\mathfrak{L}_{focal}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}-\alpha_{t}^{m,j}(1-p_{t}^{m,j})^{\gamma}\text{log}(p_{t}^{m,j})$ (8)
-    in which:
-    $\alpha_{t}^{m,j}=\begin{cases}\alpha&\text{if }y^{m}_{j}=1\\ 1-\alpha&\text{otherwise}\end{cases}$ (9)
-    $p_{t}^{m,j}=\begin{cases}\hat{y}^{m}_{j}&\text{if }y^{m}_{j}=1\\ 1-\hat{y}^{m}_{j}&\text{otherwise}\end{cases}$ (10)
-    while $\alpha$ and $\gamma$ are hyperparameters.
-  2. 2.
-    The DIoU loss [79] to optimize the regression objective by maximizing the overlap between predicted and ground truth fake time intervals:
-    $\mathfrak{L}_{DIoU}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\Bigg{(}1.0-\text{IoU}(\hat{\mathbf{b}}^{m}_{j},\mathbf{b}^{m}_{j})+\\ \frac{\rho^{2}(\hat{b}^{m}_{j},b^{m}_{j})}{\kappa^{2}}\Bigg{)}\cdot\mathds{1}_{m,j}$ (11)
-    where $j$ is frame index, $m$ is modality index, IoU denotes intersection over union, $\mathbf{b}^{m}_{j}=[s^{m}_{j},e^{m}_{j}]$ denotes a time interval, $b^{m}_{j}=0.5\cdot(e^{m}_{j}+s^{m}_{j})$ denotes the interval’s center, $\rho$ denotes the Euclidean distance, $\kappa$ denotes the length of the smallest enclosing interval covering both predicted and ground truth intervals, and $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake.
-  3. 3.
-    The smooth L1 loss [23] to minimize the distance between predicted and actual boundaries of fake time intervals:
-    $\mathfrak{L}_{1}^{smooth}=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\frac{1}{2}\bigg{(}h(s^{m}_{j}-\hat{s}^{m}_{j})\\ +h(e^{m}_{j}-\hat{e}^{m}_{j})\bigg{)}\cdot\mathds{1}_{m,j}$ (12)
-    in which $\mathds{1}_{m,j}$ is an indicator function that denotes if frame $j$ of modality $m$ is fake, and $h$ is defined as:
-    $h(x)=\begin{cases}0.5x^{2}&\text{if }\lvert x\rvert<1\\ \lvert x\rvert-0.5&\text{otherwise}\end{cases}$ (13)
-We combine the three loss functions with addition divided by the total number of fake visual and audio frames $p=\sum_{m\in\{v,a\}}\sum_{j=1}^{f}\mathds{1}_{m,j}$:
-$\mathfrak{L}_{TFL}=\big{(}\mathfrak{L}_{focal}+\mathfrak{L}_{DIoU}+\mathfrak{L}_{1}^{smooth}\big{)}/p$ (14)
-We finally average both $\mathfrak{L}_{DFD}$ and $\mathfrak{L}_{TFL}$ across batch samples during training.
-Appendix B Dataset class and split sizes
-Table 9 provides details on dataset class and split sizes.
-Although, FakeAVCeleb paper [34] reports 10,000 FVFA and 9,000 FVRA samples, the download from the corresponding homepage https://sites.google.com/view/fakeavcelebdash-lab contains 10,835 and 9,709 respectively, resulting in 21.5K total number of samples instead of the commonly reported 20,000 size for this dataset.
-Documentation therein aslo claims “Since we apply the manual screening process on synthesized videos, the final video count is more than 20,000.”.
-Test metadata of AV-Deepfake1M have not been released but evaluation is enabled through Codabench.
-FakeAVCeleb LAV-DF AV-Deepfake1M
-train val test total train val test total train val test total
-FVFA 8,678 1,052 1,105 10,835 19,090 7,701 6,369 33,160 186,344 14,515 - -
-FVRA 7,760 975 974 9,709 19,271 7,820 6,452 33,543 186,597 14,304 - -
-RVFA 399 49 52 500 19,088 7,709 6,373 33,170 186,573 14,286 - -
-RVRA 17,766 1,535 1,199 20,500 21,254 8,271 6,906 36,431 186,666 14,235 - -
-total 34,603 3,611 3,330 41,544 78,703 31,501 26,100 136,304 746,180 57,340 343,240 1,146,760
-Table 9: Dataset class and split sizes. F: fake, R: real, V: video, A: audio.
-Appendix C Generalization on FakeAVCeleb
-Table 10 provides a more detailed generalization evaluation on the FakeAVCeleb dataset with scores per forgery type, indicating the effectiveness of training on AV-Deepfake1M especially in some cases such as FVFA-WL and FVFA-FS reaching 95.6 AUC and 94.5 AUC, respectively.
-Training on LAV-DF also results in good generalization especially in FVFA-FS case with 90.0 AUC.
 Category
@@
 AP AUC AP AUC AP AUC AP AUC AP AUC AP AUC
-FakeAVCeleb [34] 84.5 97.1 99.0 99.9 99.7 99.9 99.5 99.9 99.6 99.9 99.4 99.9
-LAV-DF [10] 40.2 74.3 77.1 87.2 78.2 87.8 72.1 90.0 70.0 86.9 74.4 88.0
-AV-Deepfake1M [8] 18.8 80.8 63.5 80.3 85.6 95.6 74.2 94.5 75.9 93.8 74.8 91.1
+FakeAVCeleb [36] 66.4 51.6 100. 99.8 100. 100. 100. 100. 100. 100. 100. 99.9
+LAV-DF [10] 81.3 76.8 98.8 90.8 98.8 91.6 98.2 91.8 98.3 90.4 98.5 91.2
+AV-Deepfake1M [8] 93.9 95.5 98.2 88.6 99.7 98.1 99.4 97.8 99.4 97.5 99.2 95.5
 Table 14: Detailed generalization performance for each forgery type of the FakeAVCeleb dataset.
-Appendix D Ablation Analysis (further experiments)
-Figures 4(a) and 4(b) indicate little to no sensitivity to the focal loss’ hyperparameter $\alpha$.
-Figures 4(c) and 4(d) indicate that the use of a learning rate scheduler is beneficial however the type of scheduler is not significantly affecting the results.
-(a) Focal loss hyperparmeter $\alpha$ (LAV-DF)
-(b) Focal loss hyperparmeter $\alpha$ (AV-Deepfake1M)
-(c) Learning rate scheduler (LAV-DF)
-(d) Learning rate scheduler (AV-Deepfake1M)
-Figure 4: Ablation analysis on Temporal Forgery Localization task.
-Figure 5 indicates that on DFD window size $q=15$ is again optimal, but performance is not increased with feature pyramids nor with different learning rate schedules.
-Also, medium-sized models achieve maximum performance.
-(a) Window size $q$ (FakeAVCeleb)
-(b) Window size $q$ (LAV-DF)
-(c) Window size $q$ (AV-Deepfake1M)
-(d) Feature pyramid use (FakeAVCeleb)
-(e) Feature pyramid use (LAV-DF)
-(f) Feature pyramid use (AV-Deepfake1M)
-(g) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (FakeAVCeleb)
-(h) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (LAV-DF)
-(i) Transformer $\mathcal{T}$ size $(d,r,d\cdot u)$ (AV-Deepfake1M)
-(j) Learning rate scheduler (FakeAVCeleb)
-(k) Learning rate scheduler (LAV-DF)
-(l) Learning rate scheduler (AV-Deepfake1M)
-Figure 5: Ablation and hyperparameter tuning analysis on Deepfake Detection task.
+Table 15: Real world analysis dataset details.
+Class URL Language Video-Duration Processing-Time Coverage c
+real #1 https://www.youtube.com/watch?v=bej61P8YaXo Russian 00:00:47 0:01:31 0.198
+real #2 https://www.youtube.com/watch?v=wgx5neIx6qc Russian 00:01:16 0:01:03 0.059
+real #3 - Russian 00:02:00 0:01:36 0.177
+real #4 https://www.youtube.com/watch?v=_4PKe8WGCPg English 00:14:27 0:22:50 0.105
+real #5 https://www.youtube.com/watch?v=MBFnZhUq-Es Ukrainian 00:01:04 0:00:44 0.119
+real #6 https://www.instagram.com/zelenskyy_official/p/CaaFzibgLES/ Russian 00:00:32 0:00:27 0.063
+real #7 https://www.youtube.com/watch?v=NrDtNOiP65U Russian 00:01:08 0:00:57 0.081
+real #8 https://www.instagram.com/zelenskyy_official/p/CZKXF33KqEl/ Ukrainian 00:04:05 0:03:18 0.041
+real #9 https://www.youtube.com/watch?v=Cz2mFOJLaRI Russian 00:01:34 0:03:10 0.118
+real #10 https://www.youtube.com/watch?v=CInwveNLUmA English 00:01:34 0:02:53 0.073
+real #11 https://www.instagram.com/zelenskyy_official/p/CaBw7eNIDKp/ Ukrainian 00:01:07 0:00:55 0.028
+real #12 https://www.youtube.com/watch?v=GdxofSvTYUI English 00:03:05 0:05:40 0.031
+real #13 https://www.youtube.com/watch?v=4WpjgtK2ryo Russian 00:01:00 0:01:51 0.051
+real #14 - Italian 00:01:03 0:00:51 0.295
+real #15 https://www.youtube.com/watch?v=qPVVAy9df_Y Russian 00:11:13 0:09:27 0.415
+real #16 https://www.youtube.com/watch?v=s9s-4b5hYpw Russian 00:01:06 0:00:59 0.094
+real #17 - Italian 00:01:13 0:00:59 0.061
+real #18 https://www.youtube.com/watch?v=xVwU-YMgHvo English 00:07:01 0:10:50 0.197
+real #19 https://www.youtube.com/watch?v=rWi3T9okyvA Russian 00:02:19 0:02:23 0.053
+real #20 https://www.instagram.com/zelenskyy_official/p/CVvedjElHwX/ Russian 00:08:29 0:07:00 0.159
+real #21 https://www.youtube.com/watch?v=hHV4DEvattM Russian 00:00:56 0:01:00 0.156
+real #22 https://www.youtube.com/watch?v=HU2wY17Umzg English 00:01:03 0:02:49 0.204
+real #23 - Italian 00:02:00 0:01:24 0.025
+real #24 https://www.youtube.com/watch?v=taykmAyaw-Q Russian 00:02:19 0:01:35 0.165
+real #25 https://www.youtube.com/watch?v=Qc5hwRAXVXs Russian 00:03:54 0:04:03 0.1
+real #26 - Russian 00:02:30 0:01:42 0.343
+real #27 https://www.youtube.com/watch?v=zdH0FhnBWwY Russian 00:00:51 0:01:43 0.0
+real #28 https://www.instagram.com/zelenskyy_official/p/CakB_4yAUsa/ Ukrainian 00:06:33 0:05:21 0.611
+real #29 https://www.youtube.com/watch?v=75g7cA-_SEI Russian 00:02:12 0:04:47 0.095
+real #30 https://www.youtube.com/watch?v=Hl9Sp5xJxwk German 00:03:34 0:02:30 0.163
+real #31 https://www.youtube.com/watch?v=caxMBk1__-Y English 00:03:41 0:02:52 0.105
+real #32 - Italian 00:02:00 0:01:26 0.0
+real #33 https://www.youtube.com/watch?v=hQ58Yv6kP44 Russian 00:30:46 0:37:10 0.391
+real #34 https://www.youtube.com/watch?v=nDzX13RHTXI Arabic 00:46:26 0:53:14 0.094
+real #35 https://www.instagram.com/zelenskyy_official/p/CUHvAc9AFoX/ Ukrainian 00:01:42 0:01:20 0.13
+real #36 https://www.youtube.com/watch?v=7WUyEPkGnAc Russian 00:00:37 0:00:38 0.104
+real #37 https://www.youtube.com/watch?v=ujCFOPfqJbQ Russian 00:08:13 0:08:51 0.116
+real #38 https://www.youtube.com/watch?v=BEh1Fr-BaN8 English 00:01:46 0:01:17 0.031
+real #39 https://www.instagram.com/zelenskyy_official/p/CaklNXjAYz3/ Russian 00:14:02 0:12:36 0.085
+real #40 https://www.youtube.com/watch?v=_g_AP4hhGQM Russian 00:02:04 0:02:08 0.025
+real #41 https://www.youtube.com/watch?v=DPIR4EmaDDE Russian 00:01:16 0:01:00 0.033
+real #42 https://www.instagram.com/zelenskyy_official/p/CaokxUWARj7/ Ukrainian 00:08:25 0:13:16 0.066
+real #43 https://www.youtube.com/watch?v=rZ2J4b-dkUs English 00:00:29 0:00:58 0.535
+real #44 https://www.instagram.com/zelenskyy_official/p/CaorkbdgSv4/ Ukrainian 00:07:04 0:11:44 0.05
+real #45 https://www.instagram.com/zelenskyy_official/p/Cab8euXgtyF/ Russian 00:05:16 0:04:22 0.046
+real #46 https://www.youtube.com/watch?v=2Vmx47wsOBQ Russian 00:00:53 0:00:36 0.034
+real #47 - Italian 00:02:00 0:01:28 0.024
+real #48 - Italian 00:02:00 0:01:34 0.107
+real #49 https://www.instagram.com/zelenskyy_official/p/CaUvEpaAKWj/ Ukrainian 00:00:41 0:01:11 0.026
+real #50 https://www.youtube.com/watch?v=cxKn2sv2cgQ Ukrainian 00:05:34 0:10:49 0.048
+real #51 https://www.youtube.com/watch?v=zHdj1HW4cQw Russian 00:02:06 0:04:59 0.17
+real #52 - Russian 00:02:30 0:01:42 0.268
+real #53 https://www.youtube.com/watch?v=36_Vj3N3oIk Ukrainian 00:06:02 0:13:33 0.035
+real #54 https://www.youtube.com/watch?v=prfaWHQoxVg Russian 00:01:25 0:01:17 0.037
+real #55 https://www.youtube.com/watch?v=M0iV5vIABX0 English 00:01:36 0:01:20 0.076
+real #56 https://www.instagram.com/zelenskyy_official/p/CWh0pPuFOTA/ Ukrainian 00:07:10 0:13:56 0.09
+real #57 https://www.youtube.com/watch?v=hLCGKRF3YAI Russian 00:01:47 0:01:45 0.162
+real #58 - Russian 00:02:24 0:04:14 0.349
+real #59 - Italian 00:01:00 0:00:51 0.378
+real #60 https://www.youtube.com/watch?v=1CnyqLogH0Y Russian 00:01:40 0:01:29 0.127
+real #61 https://www.youtube.com/watch?v=WhqVwPYOY7M Ukrainian 00:03:45 0:03:00 0.14
+real #62 - Russian 00:02:27 0:04:28 0.056
+real #63 https://www.youtube.com/watch?v=TFl2nfPoxVk Russian 00:00:41 0:00:35 0.049
+real #64 https://www.instagram.com/zelenskyy_official/p/CZEnvm0KdQv/ Russian 00:01:06 0:00:55 0.025
+real #65 https://www.youtube.com/watch?v=2ze4gImZADM Russian 00:01:46 0:01:36 0.03
+real #66 https://www.youtube.com/watch?v=TccwMWVtmj0 English 00:08:02 0:07:17 0.03
+real #67 https://www.youtube.com/watch?v=J55cSx8Rz_g Russian 00:00:34 0:01:17 0.144
+real #68 https://www.instagram.com/zelenskyy_official/p/Cal8YaIFOqM/ Ukrainian 00:08:04 0:06:47 0.048
+real #69 - Russian 00:01:18 0:02:16 0.286
+real #70 - Russian 00:02:09 0:01:26 0.155
+real #71 https://www.instagram.com/zelenskyy_official/p/CYmeFavK7wC/ English 00:02:14 0:04:04 0.032
+real #72 https://www.instagram.com/zelenskyy_official/p/CaKjFkXAhts/ Ukrainian 00:19:01 0:15:29 0.336
+real #73 https://www.youtube.com/watch?v=ndNyE7f0NVs Russian 00:02:28 0:01:42 0.0
+real #74 https://www.youtube.com/watch?v=iCOwqJ8kLCM Russian 00:00:58 0:00:47 0.051
+real #75 https://www.youtube.com/watch?v=NdUkaFnunKw Russian 00:01:23 0:01:08 0.04
+fake #1 https://twitter.com/i/status/1174316798700924929 English 00:02:07 0:01:58 0.339
+fake #2 https://www.threads.net/@cj_wirenut/post/C8atfbOtQ4U English 00:00:23 0:00:19 0.342
+fake #3 https://www.dropbox.com/s/o7ndovcsycfchl3/HaoLi-UNIDIR3.mp4 English 00:00:59 0:00:47 0.415
+fake #4 https://twitter.com/reface_app/status/1352274324040908801 French 00:00:15 0:00:11 0.0
+fake #5 https://twitter.com/kaikertv/status/1284053913516179457 French 00:00:32 0:00:31 0.0
+fake #6 https://twitter.com/PhilEhr/status/1311667726742560769 English 00:00:59 0:00:49 0.119
+fake #7 https://twitter.com/facemagic_app/status/1404091187728896000 Hakha-Chin 00:00:23 0:00:18 0.0
+fake #8 https://twitter.com/ThiudaDeepfakes/status/1429455322389827592 Breton 00:00:13 0:00:18 0.849
+fake #9 https://www.instagram.com/p/C3z9bv_u2KH/ English 00:02:17 0:01:52 0.336
+fake #10 https://www.dropbox.com/s/ej7swf300aynlsz/HaoLi-UNIDIR7.mp4 English 00:00:59 0:00:48 0.426
+fake #11 https://twitter.com/dw_sports/status/1405602671525040142 French 00:00:44 0:00:57 0.0
+fake #12 https://twitter.com/facemagic_app/status/1426257244728004611 English 00:00:21 0:00:18 0.0
+fake #13 https://twitter.com/larrykim/status/1228979054579073024 French 00:01:26 0:01:12 0.22
+fake #14 https://twitter.com/EddiePozos_/status/1365420001327337474 Persian 00:00:11 0:00:10 0.603
+fake #15 https://twitter.com/sumitsaurabh/status/1324424493717049344 Tamil 00:00:09 0:00:09 0.643
+fake #16 https://twitter.com/CestCanteloup/status/1315576171803938817 French 00:00:41 0:00:35 0.041
+fake #17 https://twitter.com/welcomeai/status/1090256517666885632 Esperanto 00:00:55 0:00:38 0.4
+fake #18 https://www.facebook.com/plugins/video.php?height=476&href=https%3A%2F%2Fwww.facebook.com%2F61551008896469%2Fvideos%2F878366723849002%2F&show_text=true&width=476&t=0 English 00:01:49 0:01:36 0.279
+fake #19 https://www.youtube.com/watch?v=9ccd3LMNMl8 English 00:33:15 0:24:34 0.21
+fake #20 https://twitter.com/facemagic_app/status/1419311479065518083 Persian 00:00:10 0:00:10 0.19
+fake #21 https://fb.watch/vEHrmibPaG/ English 00:01:18 0:01:08 0.306
+fake #22 https://twitter.com/facemagic_app/status/1431270277506166796 English 00:00:11 0:00:11 0.233
+fake #23 https://packaged-media.redd.it/e4bntb4wn74a1/pb/m2-res_720p.mp4?m=DASHPlaylist.mpd&v=1&e=1743004800&s=c47b62e83cdf59e76e9a1e26a302a02091577c5c English 00:00:16 0:00:16 0.546
+fake #24 https://twitter.com/facemagic_app/status/1408439846474072064 Persian 00:00:12 0:00:11 0.0
+fake #25 https://twitter.com/facemagic_app/status/1421485812609323011 Spanish 00:00:12 0:00:12 0.223
+fake #26 https://twitter.com/facemagic_app/status/1430545513955221508 Hakha-Chin 00:00:11 0:00:12 0.463
+fake #27 https://www.facebook.com/plugins/video.php?height=476&href=https%3A%2F%2Fwww.facebook.com%2F61552447227376%2Fvideos%2F870526964468908%2F&show_text=false&width=476&t=0 English 00:01:42 0:01:30 0.159
+fake #28 https://twitter.com/hytel/status/1425631238208081924 English 00:00:43 0:00:38 0.275
+fake #29 https://twitter.com/Happy_Finish/status/1091007206944919559 English 00:01:16 0:01:08 0.264
+fake #30 https://twitter.com/FutureAdvocacy/status/1194148111121354752 Welsh 00:01:11 0:00:59 0.045
+fake #31 https://twitter.com/facemagic_app/status/1404453586382188550 English 00:00:14 0:00:14 0.0
+fake #32 https://twitter.com/facemagic_app/status/1426559247043674112 Persian 00:00:15 0:00:13 0.0
+fake #33 https://twitter.com/facemagic_app/status/1428371186468098064 Breton 00:00:44 0:00:36 0.427
+fake #34 https://twitter.com/brianmonarch/status/1307037446056677382 English 00:00:59 0:00:54 0.446
+fake #35 https://www.bitchute.com/video/8ZGoBmGmKtgi English 00:00:37 0:00:32 0.369
+fake #36 https://twitter.com/facemagic_app/status/1424384912723824641 French 00:00:17 0:00:14 0.0
+fake #37 https://www.bitchute.com/video/ChybNNbjEEOd English 00:02:16 0:01:55 0.114
+fake #38 https://twitter.com/CestCanteloup/status/1277896113618669569 French 00:00:51 0:00:39 0.153
+fake #39 https://www.dropbox.com/s/5ll22su3gzc4qyn/HaoLi-UNIDIR4.mp4 English 00:00:59 0:00:47 0.201
+fake #40 https://twitter.com/facemagic_app/status/1406627903249518596 English 00:00:23 0:00:18 0.683
+fake #41 https://www.dropbox.com/s/b9o21z834kwi59s/HaoLi-UNIDIR1.mp4 English 00:00:59 0:00:46 0.504
+fake #42 https://twitter.com/insanedio/status/1308620241732157440 Spanish 00:00:44 0:00:35 0.391
+fake #43 https://www.youtube.com/watch?v=J7WqjrtJHkk Welsh 00:00:50 0:00:42 0.07
+fake #44 https://www.facebook.com/100092555835479/posts/1389999691891907/ Romanian 00:00:15 0:00:14 0.115
+fake #45 https://twitter.com/facemagic_app/status/1410251795348271126 English 00:00:19 0:00:16 0.058
+fake #46 https://www.facebook.com/100095605808597/posts/3495941347387764/ Romanian 00:01:04 0:00:45 0.79
+fake #47 https://twitter.com/stephaniemain2/status/1315238754148212736 Esperanto 00:00:15 0:00:10 0.322
+fake #48 https://twitter.com/Seekthetruth101/status/1308135272706695169 English 00:02:09 0:01:48 0.174
+fake #49 https://twitter.com/facemagic_app/status/1416774762441830406 English 00:00:16 0:00:13 0.594
+fake #50 https://twitter.com/facemagic_app/status/1426921628462567424 Dhivehi 00:00:13 0:00:13 0.752
+fake #51 https://twitter.com/EvanKirstel/status/1268984667153276928 English 00:01:01 0:00:51 0.174
+fake #52 https://twitter.com/drfakenstein/status/1259838123468632065 English 00:00:57 0:00:47 0.431
+fake #53 https://twitter.com/facemagic_app/status/1429095960093790210 English 00:00:24 0:00:20 0.259
+fake #54 https://www.bitchute.com/video/sq78wmU60uHY German 00:00:44 0:01:08 0.365
+fake #55 https://twitter.com/KRMTRoySuryo2/status/1327286419857915910 Spanish 00:00:50 0:00:37 0.801
+fake #56 https://media2.ellinikahoaxes.gr/uploads/2024/10/video-isxurismou-deepfake.mp4 English 00:00:09 0:00:09 0.04
+fake #57 https://www.facebook.com/EtzrodtJaimeeShuler/posts/7939864682747869/ English 00:02:03 0:01:55 0.322
+fake #58 https://twitter.com/sp_a/status/998089909369016325 English 00:00:54 0:00:39 0.269
+fake #59 https://twitter.com/drfakenstein/status/1171638420822736897 Chinese-Taiwan 00:00:59 0:00:49 0.017
+fake #60 https://twitter.com/facemagic_app/status/1420761028468649985 English 00:00:16 0:00:15 0.032
+fake #61 https://www.facebook.com/plugins/video.php?height=476&href=https%3A%2F%2Fwww.facebook.com%2F100089533433509%2Fvideos%2F897022265405097%2F&show_text=false&width=476&t=0 English 00:01:40 0:01:27 0.292
+fake #62 https://twitter.com/hwingo/status/1090809957342154752 English 00:01:35 0:01:09 0.089
+fake #63 https://twitter.com/facemagic_app/status/1426196858418843655 Chinese-Taiwan 00:00:22 0:00:17 0.253
+fake #64 https://www.facebook.com/plugins/video.php?height=476&href=https%3A%2F%2Fwww.facebook.com%2F61550777281885%2Fvideos%2F1992291657812664%2F&show_text=false&width=476&t=0 English 00:01:40 0:01:26 0.323
+fake #65 https://web.archive.org/web/20240816152954if_/https://video.xx.fbcdn.net/v/t42.1790-2/453528538_3676577589299215_6162862663998579296_n.mp4?_nc_cat=106&ccb=1-7&_nc_sid=55d0d3&efg=eyJ2ZW5jb2RlX3RhZyI6InN2ZV9zZCIsInZpZGVvX2lkIjo1MDc0Nzc3NTQ5OTkxNjF9&_nc_ohc=fI8yJJw2qCkQ7kNvgG5DX0W&_nc_ht=scontent.fprg1-1.fna&oh=00_AYD-yhNK-2jlWPV8DVBgUL2KpMUOMiMbue1ss52pCISILg&oe=66C548DA Czech 00:01:03 0:00:50 0.661
+fake #66 https://twitter.com/facemagic_app/status/1401916866038169601 English 00:00:25 0:00:20 0.809
+fake #67 https://video.twimg.com/ext_tw_video/1407579338531291138/pu/vid/320x568/nNPZ0rn3C9kFsr_P.mp4 Russian 00:00:19 0:00:17 0.242
+fake #68 https://twitter.com/facemagic_app/status/1409889406404726784 English 00:00:16 0:00:15 0.0
+fake #69 https://twitter.com/adrlenard/status/1172567432239751168 English 00:00:21 0:00:18 0.264
+fake #70 https://twitter.com/facemagic_app/status/1428008798317387777 Esperanto 00:00:27 0:00:19 0.0
+fake #71 https://twitter.com/i/status/1095597694712799232 French 00:00:16 0:00:14 0.0
+fake #72 https://twitter.com/PapillonFlacko/status/1282885977518809089 Kabyle 00:02:18 0:01:40 0.326
+fake #73 https://twitter.com/facemagic_app/status/1417861927808606212 French 00:00:10 0:00:11 0.72
+fake #74 https://twitter.com/facemagic_app/status/1406265519234224132 Hakha-Chin 00:00:19 0:00:16 0.424
+fake #75 https://www.bitchute.com/video/bvJsG8fyeCTt English 00:00:26 0:00:23 0.192
+fake #76 - English 00:00:27 0:00:22 0.355
+fake #77 https://www.tiktok.com/@jbviet/video/6994005244309032197 French 00:00:18 0:00:16 0.136
+fake #78 https://twitter.com/facemagic_app/status/1424747301277364229 Romansh-Sursilvan 00:00:18 0:00:14 0.613
+fake #79 https://www.facebook.com/Jmvgames/videos/821549489686229 English 00:01:16 0:01:05 0.447
+fake #80 https://twitter.com/facemagic_app/status/1430183118808387585 Welsh 00:00:11 0:00:12 0.038
+fake #81 https://twitter.com/CestCanteloup/status/1276794899355049984 French 00:00:46 0:00:37 0.107
+fake #82 https://twitter.com/Nosetwittears/status/1425532656532475919 English 00:00:17 0:00:14 0.055
+fake #83 https://x.com/TheBabylonBee/status/1836518555648692570 English 00:01:55 0:01:36 0.151
+fake #84 https://twitter.com/GACrewFans/status/1308559503462297600 English 00:00:14 0:00:11 0.0
+fake #85 https://www.facebook.com/plugins/video.php?height=476&href=https%3A%2F%2Fwww.facebook.com%2FNwadshairnnails%2Fvideos%2F203824829410862%2F&show_text=false&width=476&t=0 English 00:01:42 0:01:27 0.303
+fake #86 https://twitter.com/RT_com/status/1308410711027089411 Dhivehi 00:02:02 0:01:42 0.191
+fake #87 https://twitter.com/Knuckle_HeadTV/status/1426780954056658947 Romansh-Sursilvan 00:01:00 0:00:50 0.216
+fake #88 https://twitter.com/GeorgeTakei/status/1269459672714694661 English 00:01:22 0:02:02 0.025
+fake #89 https://www.dropbox.com/s/d0qfadmo05ckdoj/HaoLi-UNIDIR8.mp4 English 00:00:59 0:00:50 0.508
+fake #90 https://twitter.com/nuweverse/status/1320405519438405634 English 00:00:10 0:00:10 0.35
+fake #91 https://twitter.com/NuritBen/status/1284084875897778177 Welsh 00:01:09 0:01:02 0.107
+fake #92 https://media2.ellinikahoaxes.gr/uploads/2024/06/447088377_986072086506420_2750102588021047120_n.mp4 Greek 00:03:57 0:03:08 0.671
+fake #93 https://twitter.com/simerazzi/status/1333843243516223488 Kabyle 00:00:15 0:00:12 0.907
+fake #94 https://twitter.com/facemagic_app/status/1411701337981661192 Hakha-Chin 00:00:08 0:00:09 0.236
+fake #95 https://twitter.com/facemagic_app/status/1427284018979090433 English 00:00:33 0:00:26 0.436
+fake #96 https://twitter.com/callmeuschi/status/1377599278252707845?s=21 English 00:00:17 0:00:14 0.231
+fake #97 https://twitter.com/facemagic_app/status/1427646404344700928 English 00:00:16 0:00:15 0.033
+fake #98 https://twitter.com/facemagic_app/status/1411338949977591812 English 00:00:18 0:00:15 0.303
+fake #99 https://twitter.com/facemagic_app/status/1430183114337361931 English 00:00:28 0:00:21 0.131
+fake #100 - English 00:00:28 0:00:26 0.079
+fake #101 https://www.dailymotion.com/video/x7mf9hf English 00:03:02 0:02:11 0.122
+fake #102 https://www.dropbox.com/s/6v58miw1o7c015z/HaoLi-UNIDIR2.mp4 English 00:00:59 0:00:49 0.491
+fake #103 https://www.facebook.com/100001934236498/videos/594209412579621/ English 00:01:39 0:01:23 0.252
+fake #104 https://www.dropbox.com/s/8n7o5ajkllw4wgm/HaoLi-UNIDIR6.mp4 English 00:00:59 0:00:47 0.444
+fake #105 https://web.archive.org/web/20240821053915if_/https://video-lga3-2.xx.fbcdn.net/o1/v/t2/f2/m69/An862HN9EEl8l00qEY6Y09SfPqP08f9VqMtlf4F9ZkQ8cY4H5jrsoEyvCLx4MDKQy0EqvFeuKwZ_1zS4s4fmXTNa.mp4?efg=eyJ2ZW5jb2RlX3RhZyI6Im9lcF9oZCJ9&_nc_ht=video-lga3-2.xx.fbcdn.net&_nc_cat=111&strext=1&vs=fb8e3bd95d6a8122&_nc_vs=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%3D&ccb=9-4&oh=00_AYByVnpiEhIPvm2WhRSvuqB-gWs_wYumgmOnAvtFGL4NfA&oe=66C748CE&_nc_sid=1d576d&_nc_rid=779600729108004&_nc_store_type= Czech 00:00:53 0:00:46 0.242
+fake #106 https://media2.ellinikahoaxes.gr/uploads/2024/06/447071239_3323218401145262_6755718274625284919_n.mp4 Greek 00:03:58 0:03:15 0.673
+fake #107 https://twitter.com/facemagic_app/status/1415325224468369413 English 00:00:12 0:00:12 0.542
+fake #108 https://twitter.com/facemagic_app/status/1431995054181208066 Hakha-Chin 00:00:12 0:00:13 0.0
+fake #109 https://twitter.com/DBTagsen_Ofc/status/1312217715050979328 English 00:00:16 0:00:16 0.0
+fake #110 https://twitter.com/facemagic_app/status/1430907889305194496 English 00:00:10 0:00:11 0.894
+fake #111 https://twitter.com/SuperHeroAcade/status/1308803743421468677 English 00:00:14 0:00:14 0.329
+fake #112 https://twitter.com/DrDMemes1/status/1306246535160754177 English 00:01:44 0:01:37 0.186
+fake #113 https://twitter.com/facemagic_app/status/1406990302083272704 Dhivehi 00:00:15 0:00:14 0.0
+fake #114 https://www.tiktok.com/@jbviet/video/6995159085595331846 Chinese-China 00:00:11 0:00:11 0.12
+fake #115 https://twitter.com/AllanXia/status/1168521996168097793 Persian 00:00:26 0:00:35 0.373
+fake #116 https://twitter.com/facemagic_app/status/1414600441049010182 English 00:00:21 0:00:16 0.247
+fake #117 https://twitter.com/wanmafiq/status/1330742746869420032%****␣appendix.tex␣Line␣800␣**** Spanish 00:00:09 0:00:08 0.0
+fake #118 https://twitter.com/versl9/status/1281132787089711104 Russian 00:00:42 0:00:36 0.446
+fake #119 https://twitter.com/firejapan/status/1305133253225922560 Breton 00:00:07 0:00:08 0.542
+fake #120 https://twitter.com/stephaniemain2/status/1326246214078238720 Chinese-Taiwan 00:00:11 0:00:10 0.542
+fake #121 https://twitter.com/drfakenstein/status/1171182601563820038 English 00:00:27 0:00:25 0.602
+fake #122 https://twitter.com/rschu/status/1182372560236888065 English 00:02:19 0:02:08 0.218
+fake #123 https://www.facebook.com/MISTERKOCTWO/videos/949168468846018/ French 00:01:00 0:00:39 0.0
+fake #124 https://twitter.com/facemagic_app/status/1432357446450749446 Spanish 00:00:12 0:00:12 0.0
+fake #125 https://twitter.com/facemagic_app/status/1421848193671712770 English 00:00:13 0:00:12 0.087
+fake #126 https://twitter.com/CestCanteloup/status/1275394966022471685 French 00:00:37 0:00:33 0.0
+fake #127 https://twitter.com/lachoseparis/status/1181168724629475328 English 00:00:32 0:00:24 0.054
+fake #128 https://twitter.com/Startup_Nerd/status/1283695456992862208 English 00:00:49 0:00:38 0.094
+fake #129 https://twitter.com/MikaelThalen/status/1311696656061009920 English 00:00:49 0:00:37 0.166
+(a) Real predicted as fake: LAV-DF/test/000020.mp4
+(b) Fake predicted as real: LAV-DF/test/000038.mp4
+Figure 16: Example videos that DiMoDif has misclassified.